### v1

# On the Hurwitz Stability of Hurwitz-Type Matrix Polynomials

Abstract
Every matrix polynomial ${\bf f}_{n}$ can be written in the form ${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2})$. The matrix polynomial ${\bf f}_{2m}$ is said to be of Hurwitz-type if the expression ${\bf g}_{2m}(z){\bf h}_{2m}^{-1}(z)$ admits a representation as a finite continued fraction with positive definite matrix coefficients. Similarly, the odd-degree matrix polynomial ${\bf f}_{2m+1}$ is of Hurwitz type if $\frac{1}{z}{\bf h}_{2m+1}(z){\bf g}_{2m+1}^{-1}(z)$ has the same property. Using orthogonal matrix polynomials on the interval $[0,+\infty)$ and their associated second-kind polynomials, we establish the coprimeness of the matrix polynomial pairs $({\bf h}_{2m},{\bf g}_{2m})$ and $({\bf h}_{2m+1},z{\bf g}_{2m+1})$. Assuming a commutativity-type condition on the coefficients of $f_{n}$, we derive an explicit form of the Bezoutian associated with Hurwitz-type matrix polynomials. Under the same assumption, we also show that Hurwitz-type matrix polynomials are indeed Hurwitz matrix polynomials. Finally, we propose an extension of the class of Hurwitz-type matrix polynomials.
AMS Subject Classification (2020): 34D20 33C45 47A56 15A24 30E05
Keywords: Hurwitz matrix polynomial, orthogonal matrix polynomial, Stieltjes transform, Bezoutian
1 Introduction
A scalar polynomial with all its roots in the left half of the complex plane is called a Hurwitz polynomial [60, 41, 58, 19, 32, 40, 46, 10, 42, 43]. The Hurwitz stability of such polynomials is also referred to as Hurwitzness—a term we adopt in this work. Let $n$ and $q$ be natural numbers. In the present work, we consider the properties and Hurwitzness of a subset of Lambda matrices [55, page 95], [33, page 228], [20], defined by
${\bf f}_{n}(z):=A_{0}z^{n}+A_{1}z^{n-1}+\ldots+A_{n},$ (1)
where $A_{j}$ is a complex $q\times q$ matrix and $z$ is a complex variable. As in [50], we refer to a matrix of the form (1) as a matrix polynomial, which is called a Hurwitz matrix polynomial if $\det{\bf f}_{n}$ is a Hurwitz polynomial.
The matrix polynomial (1) has degree $n$ if $A_{0}$ is nonzero. Throughout this paper, we assume that $\det A_{0}\neq 0$. Clearly, (1) can be represented as
${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2}),$ (2)
where
${\bf h}_{n}(z)$ :=$:=\left\{\begin{array}[]{ll}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{array}\right.$ (5)
${\bf h}_{n}(z)$ :=$:=\begin{cases}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{cases}$ (6)
and
${\bf g}_{n}(z)$ :=$:=\left\{\begin{array}[]{ll}A_{1}z^{m-1}+A_{3}z^{m-2}+\ldots+A_{2m-1},&n=2m,\\ A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m+1.\end{array}\right.$ (9)
We will refer to (5) as the ${\bf h}_{n}$ part and to (9) as the ${\bf g}_{n}$ part of the matrix polynomial ${\bf f}_{n}$.
Let us introduce the following notation. The symbol ${\mathbb{C}}^{q\times q}$ represents the set of all complex $q\times q$ matrices. For the null matrix that belongs to ${\mathbb{C}}^{p\times q}$, we will write $0_{p\times q}$. We denote by $0_{q}$ and $I_{q}$ the null and identity matrices in ${\mathbb{C}}^{q\times q}$, respectively. When the sizes of the null and identity matrices are clear from the context, we will omit their indices.
In the following definition, we recall the concept of a Hurwitz-type matrix (HTM) polynomial [8]. In [8], HTM polynomials were referred to as matrix Hurwitz-type polynomials. For $A,\,B\in{\mathbb{C}}^{q\times q}$ with $B$ invertible, set $\frac{A}{B}:=AB^{-1}$.
Definition 1.1.
The $q\times q$ matrix polynomial ${\bf f}_{n}$ in (1) is called a Hurwitz-type matrix polynomial if there exist two sequences of positive definite matrices, $({\bf c}_{k})_{k=0}^{m}$ and $({\bf d}_{k})_{k=0}^{m-1}$ (resp. $({\bf c}_{k})_{k=0}^{m}$ and $({\bf d}_{k})_{k=0}^{m}$) such that, for $n=2m$,
$\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}=$ $\cfrac{I_{q}}{z{\bf c}_{0}+\cfrac{I_{q}}{{\bf d}_{0}+\cfrac{I_{q}}{\ddots+{\bf d}_{m-2}+\cfrac{I_{q}}{+z{\bf c}_{m-1}+{\bf d}_{m-1}^{-1}}}}},$ (10)
for all $z\in\mathbb{C}$ with $\det{\bf h}_{n}(z)\neq 0$ and, for $n=2m+1$,
$\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}=$ $\cfrac{I_{q}}{z{\bf c}_{0}+\cfrac{I_{q}}{{\bf d}_{0}+\cfrac{I_{q}}{\ddots\,+z{\bf c}_{m-1}+\cfrac{I_{q}}{{\bf d}_{m-1}+z^{-1}{\bf c}_{m}^{-1}}}}}$ (11)
for all $z\in\mathbb{C}\setminus\{0\}$ with $\det{\bf g}_{n}(z)\neq 0$.
The definition of HTM polynomials in the form of a finite continued fraction is inspired by [34, Chapter XV, Theorem 15].
The rational matrix functions in (10) and (11) are special solutions to the truncated Stieltjes matrix moment (TSMM) problem; see proof of [8, Theorem 7.9]. This implies that $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ are Stieltjes transforms of a certain positive measure $\sigma$ on $[0,+\infty)$.
The TSMM problem is defined as follows: Let $(s_{j})_{j=0}^{n}$ be a finite sequence of $q\times q$ complex Hermitian matrices. The TSMM problem consists of finding the set ${\mathscr{M}}_{n}$ of all nonnegative Hermitian $q\times q$ measures $\sigma$ defined on the Borel $\sigma$–algebra ${\mathfrak{B}}\cap[0,+\infty)$ such that
$s_{j}=\int\limits_{[0,+\infty)}t^{j}\sigma(dt),\quad 0\leq j\leq n-1,\quad s_{n}=\int\limits_{[0,+\infty)}t^{n}\sigma(dt)+M,$ (12)
where $M$ is a $q\times q$ complex valued positive semi-definite matrix.
For each $\sigma\in{\mathscr{M}}_{n}$, one typically associates it with a matrix function
$s(z)=\int_{[0,+\infty)}\frac{\sigma(dt)}{t-z},$ (13)
which is defined and holomorphic in $\mathbb{C}\setminus[0,+\infty)$. The matrix function $s$ is said to be associated with the TSMM problem. See [21, 22, 23, 24, 26, 27, 17, 29], [30, 31].
In the present work, we use the following condition on the coefficients of the matrix polynomial ${\bf f}_{n}$.
Condition C. For $n=2m$ (respectively, $n=2m+1$), assume that
$A_{2j+1}^{*}A_{2k}=A_{2k}^{*}A_{2j+1},$ (14)
for all $j=0,1,\ldots,m-1$ and $k=0,1,2,\ldots,m$ (respectively, $j=0,1,\ldots,m$ and $k=0,1,2,\ldots,m$).
This condition is a consequence of [2, Remark 2.2], which outlines the conditions for the Bezoutian of a rational matrix function to admit two distinct matrix factorizations. See also Remark 2.18.
Main Results. Below are the main results of the present work:
  • a)
    We prove the coprimeness of the pair of matrix polynomials $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ (resp. $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$); see Theorem 3.5.
  • b)
    Under Condition C (14), we compute explicitly the Bezoutian of the HTM polynomial (1); see Theorem 4.4 and Corollary 4.5.
  • c)
    Under the same condition as in part b), we verify the Hurwitzness of HTM polynomials; see Theorem 5.3.
  • d)
    For a given matrix polynomial $P_{n}$ of degree $n$, which is not a Hurwitz-type matrix (HTM) polynomial, we construct a polynomial $f_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ of degree $2n$, with a suitably chosen $Q_{n-1}$, such that ${\bf f}_{2n}$ becomes an HTM polynomial. Then, by applying the result from [25], we conclude that $P_{n}$ itself is an HTM polynomial; see Theorem 5.11.
The coprimeness of matrix polynomials was previously studied in [44, 66, 45, 3]. The Hurwitzness of matrix polynomials via Bezoutians was considered in [37, 53, 54]. Note that in [67] the Hurwitzness of the HTM polynomials was examined but without leveraging the commutative-type properties of the coefficients of the HTM polynomials. Moreover, the proof of [67, Theorem 3.5], which serves as the basis for the analysis of the Hurwitzness of HTM polynomials in [67], is not valid; see Remark 5.8.
Additionally, in contrast to [67], we use the representations of the matrix polynomial ${\bf f}_{n}$ as in (2), (74) and Condition C given by (14) to explicitly compute the Bezoutian of the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$). The asterisk denotes the Hermitian adjoint of a matrix, while the overbar represents complex conjugation. In the scalar case, the use of the ${\bf h}_{n}$ part and ${\bf g}_{n}$ part of the polynomial ${\bf f}_{n}$ for studying the corresponding Bezoutian was employed by Krein and Naimark in [48] and by Krein and Akhiezer in [1].
The importance of HTM polynomials as a subset of Hurwitz matrix polynomials is that they help study the asymptotic stability of the differential equation
$I_{q}y^{(n)}(t)+A_{1}y^{(n-1)}(t)+\ldots+A_{n-1}y^{\prime}(t)+A_{m}y(t)=0,\ \ t\geq 0$ (15)
with $y(t)\in{\mathbb{R}}^{n}$; $I_{q},A_{1},\ldots,A_{n}\in\mathbb{C}^{q\times q}$; see [56, 63, 59, 38] and [47].
Our motivation is to generalize the results regarding robust stability [9] and finite-time stabilization with bounded controls [15, 16, 11, 13] for the matrix case within the framework of HTM polynomials.
Our work is organized as follows. In Section 2, we review previous results regarding HTM polynomials and their associated Markov parameters. Moreover, we mention Stieltjes nonnegative definite extendable sequences, Stieltjes orthogonal matrix polynomials relatively prime polynomial matrices, the Bezoutian, and the inertia of matrix polynomials. In Section 3, we examine the properties of HTM polynomials by exploring the relationships between the Markov parameters and the coefficients of HTM polynomials, as well as the interrelations between HTM polynomials and Markov parameters. Furthermore, we analyze the coprimeness of the matrix polynomial pairs $({\bf h}_{2m},{\bf g}_{2m})$ and $({\bf h}_{2m+1},z{\bf g}_{2m+1})$. In Section 4, we explicitly compute the Bezoutian of HTM polynomials. In Section 5, we study the Hurwitzness of HTM polynomials and we propose a procedure to derive an HTM from a given matrix polynomial.
This work is a direct continuation of the author’s research in [7] and [8], which focus on the nondegenerate truncated matricial Stieltjes problem and the HTM polynomials, respectively.
2 Preliminaries
In this section, we present the notions and results that will be used throughout this work.
2.1 HTM polynomials and associated Markov parameters
In the following remark, we reproduce the Laurent expansion of $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$, which allows us to relate the matrix polynomial (1) with the so-called Markov parameters $s_{j}$ of the polynomial ${\bf f}_{n}$.
Remark 2.1.
Let ${\bf h}_{n}$ and ${\bf g}_{n}$ be defined as in (5) and (9), and let ${\bf f}_{n}$ be as in (2). Let $|z_{0}|\geq\max\{|\lambda|:\det{\bf h}_{n}(\lambda)=0\}$ (resp. $|z_{0}|\geq\max\{|\lambda|:\det{\bf g}_{n}(\lambda)=0\}$). For $|z|>|z_{0}|$, the power series expansion of ${\bf g}_{n}/{\bf h}_{n}$ (resp. ${\bf h}_{n}/(z{\bf g}_{n}$)) in negative powers of $z$ yields
$\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ $=\frac{s_{0}}{z}-\frac{s_{1}}{z^{2}}+\ldots+(-1)^{n}\frac{s_{n}}{z^{n+1}}+\ldots\,,\ \ n=2m,$ (16)
$\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ $=\frac{s_{0}}{z}-\frac{s_{1}}{z^{2}}+\ldots+(-1)^{n}\frac{s_{n}}{z^{n+1}}+\ldots\,,\ \ n=2m+1.$ (17)
For $j\geq 0$ and by using the Markov parameters, we construct block Hankel matrices:
$H_{1,j}:=\left(\begin{array}[]{cccc}s_{0}&s_{1}&\ldots&s_{j}\\ s_{1}&s_{2}&\ldots&s_{j+1}\\ \vdots&\vdots&\vdots&\vdots\\ s_{j}&s_{j+1}&\ldots&s_{2j}\end{array}\right),H_{2,j}:=\left(\begin{array}[]{cccc}s_{1}&s_{2}&\ldots&s_{j+1}\\ s_{2}&s_{3}&\ldots&s_{j+2}\\ \vdots&\vdots&\vdots&\vdots\\ s_{j+1}&s_{j+2}&\ldots&s_{2j+1}\end{array}\right).$ (18)
2.2 The Hankel nonnegative definite extendable sequences
Consider a TSMM problem with an odd (resp. even) number of moments $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) as in (18). In [21], it is proven that the TSMM problem in the case of an odd (resp. even) number of moments has a solution if and only if $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are both positive semi-definite matrices.
Definition 2.2.
The sequence $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is called a Stieltjes positive sequence if both corresponding matrices $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are positive definite. In the case where $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is a positive Stieltjes sequence, the corresponding TSMM problem is called a TSMM problem in the nondegenerate case.
In [21] and [22], Yu. Dyukarev found the complete set of solutions to the TSMM problem in the case that $(s_{j})_{j=0}^{2m}$ and $(s_{j})_{j=0}^{2m+1}$ are Stieltjes positive sequences, respectively.
We recall the concept of Stieltjes nonnegative definite extendable sequences for the TSMM problem for $n=2m+1$ and $n=2m$; refer to [30] for $\alpha=0$. See also [6], [27] and references therein.
Definition 2.3.
Let $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) be a sequence from ${\mathbb{C}}^{q\times q}$ for which there exists some $s_{2m}\in{\mathbb{C}}^{q\times q}$ (resp. $s_{2m+1}\in{\mathbb{C}}^{q\times q}$) such that $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are Hankel nonnegative definite block matrices. The sequence $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) is called a Stieltjes nonnegative definite extendable sequence.
Next we paraphrase [30, Proposition 2.20] regarding Stieltjes nonnegative definite extendable sequences for $\alpha=0$.
Remark 2.4.
The set of Stieltjes positive sequences $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is a subset of the set of Stieltjes nonnegative definite extendable sequences $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$).
Let
$Y_{1,j}$ $:=y_{[j,2j-1]},\,\ 1\leq j\leq n,\,\quad\mbox{and}\quad Y_{2,j}:=y_{[j+1,2j]},\,\ 2\leq j\leq n,$ (19)
with
$y_{[j,k]}:=\left(s_{j},s_{j+1},\ldots,s_{k}\right)^{*},\quad 0\leq j\leq k$ (20)
and $y_{[j,k]}=0_{q}$ if $j>k$.
In the following proposition, we restate Propositions 4.13 and 4.14 from [30] for $\alpha=0$ using our notation. If $A$ and $B$ are $q\times q$ matrices, the inequality $A>B$ means that the matrix $A-B$ is positive definite.
Proposition 2.5.
a) The sequence $(s_{j})_{j=0}^{2m}$ is a positive Stieltjes sequence if and only if $(s_{j})_{j=0}^{2m-1}$ is a positive Stieltjes sequence and
$s_{2m}>Y_{1,m}^{*}H_{1,m-1}^{-1}Y_{1,m}.$ (21)
b) The sequence $(s_{j})_{j=0}^{2m+1}$ is a positive Stieltjes sequence if and only if $(s_{j})_{j=0}^{2m}$ is a positive Stieltjes sequence and
$s_{2m+1}>Y_{2,m}^{*}H_{2,m-1}^{-1}Y_{2,m}.$ (22)
Note that Proposition 2.5 indicates out the condition that matrices $s_{2m}$ and $s_{2m+1}$ must satisfy to be the extension of sequences $(s_{j})_{j=0}^{2m-1}$ and $(s_{j})_{j=0}^{2m}$, respectively.
2.3 Stieltjes orthogonal matrix polynomials
We now reproduce from [7] a number of auxiliary matrices, as well as two families of orthogonal matrix polynomials and their polynomials of the second kind. Let $R_{j}:\mathbb{C}\to\mathbb{C}^{(j+1)q\times(j+1)q}$ be given by
$R_{j}(z)$ $:=(I_{(j+1)q}-zT_{j})^{-1},\qquad j\geq 0,$ (23)
with
$T_{0}:=0_{q},\ \ T_{j}:=\left(\begin{array}[]{cc}0_{q\times jq}&0_{q}\\ I_{jq}&0_{jq\times q}\\ \end{array}\right),\qquad j\geq 1.$ (24)
Observe that for each $j\in\mathbb{N}_{0}$, the matrix–valued function $R_{j}$ can be represented via
$R_{j}(z)=\left(\begin{array}[]{cccccc}I_{q}&0_{q}&0_{q}&\ldots&0_{q}&0_{q}\\ zI_{q}&I_{q}&0_{q}&\ldots&0_{q}&0_{q}\\ z^{2}I_{q}&zI_{q}&I_{q}&\ldots&0_{q}&0_{q}\\ \vdots&\vdots&\vdots&\iddots&\vdots&\vdots\\ z^{j}I_{q}&z^{j-1}I_{q}&z^{j-2}I_{q}&\ldots&zI_{q}&I_{q}\end{array}\right).$ (25)
Let
$v_{0}:=$ $I_{q},\quad v_{j}:=\left(\begin{array}[]{c}I_{q}\\ 0_{jq\times q}\\ \end{array}\right)\ \,\mbox{for all}\ \,j\in\mathbb{N}.$ (28)
Furthermore, let $y_{[j,k]}$ be as in (20), and define
$u_{1,0}:=0_{q},\quad u_{1,j}:=\left(\begin{array}[]{c}0_{q}\\ -y_{[0,j-1]}\\ \end{array}\right),\quad u_{2,j}:=\begin{array}[]{c}-y_{[0,j]}\\ \end{array},\quad 1\leq j\leq n.$ (32)
Let $\widehat{H}_{1,j}$ (resp. $\widehat{H}_{2,j}$) denote the Schur complement of the block Hankel matrices $H_{1,j-1}$ in $H_{1,j}$ (resp.  $H_{2,j-1}$ in $H_{2,j}$) as seen in [57], [7, Equalities (2.5) and (2.6)]:
$\widehat{H}_{1,0}:=s_{0},\ \ \widehat{H}_{1,j}:=$ $s_{2j}-Y_{1,j}^{*}H_{1,j-1}^{-1}Y_{1,j},\ \ j\geq 1,$ (33)
$\widehat{H}_{2,0}:=s_{1},\ \ \widehat{H}_{2,j}:=$ $s_{2j+1}-Y_{2,j}^{*}H_{2,j-1}^{-1}Y_{2,j},\ \ j\geq 1.$ (34)
These matrices are Hermitian positive definite matrices, and so are $H_{1,j}$ and $H_{2,j}$.
The positive definite matrices in (35) to (37) of the following definition were introduced in [23] by Yu. Dyukarev.
Definition 2.6.
Let $(s_{j})_{j=0}^{2n}$ (resp. $(s_{j})_{j=0}^{2n+1}$) be a Stieltjes positive sequence. Furthermore, let $v_{j}$ and $u_{2,j}$ be as in (28) and (32), respectively. The $q\times q$ matrices
$M_{0}:=$ $s_{0}^{-1},\quad L_{0}:=s_{0}s_{1}^{-1}s_{0},$ (35)
$M_{j}:=$ $v_{j}^{*}H_{1,j}^{-1}v_{j}-v_{j-1}^{*}H_{1,j-1}^{-1}v_{j-1},$ (36)
$L_{j}:=$ $u_{2,j}^{*}H_{2,j}^{-1}u_{2,j}-u_{2,j-1}^{*}H_{2,j-1}^{-1}u_{2,j-1}$ (37)
are called the Dyukarev-Stieltjes matrix parameters of the TSMM problem.
T.J. Stieltjes’s renowned 1894 work “Recherches sur les fractions continues” [64] introduced the scalar version of the Dyukarev-Stieltjes parameters. These parameters, known as the Stieltjes parameters ([35], [49]), are given by $l_{j}:=\frac{\Delta_{j}^{2}}{\Delta_{j}^{(1)}\Delta_{j-1}^{(1)}}$, $m_{j}:=\frac{[\Delta_{j-1}^{(1)}]^{2}}{\Delta_{j}\Delta_{j-1}}$ and $j\geq 0$ with $\Delta_{j}:=\det(s_{i+k})_{i,k=0}^{j}$, $\Delta_{j}^{(1)}:=\det(s_{i+k+1})_{i,k=0}^{j}$, ($\Delta_{-1}:=\Delta_{-1}^{(1)}=1$).
Now we partially reproduce left orthogonal matrix polynomials (OMP) on $[0,+\infty)$ [25].
Definition 2.7.
Let $\sigma$ be a $q\times q$ nonnegenative measure on $[0,+\infty)$. A sequence $(P_{j})_{j=0}^{\infty}$ of $q\times q$ matrix-valued polynomials is called orthogonal with respect to $\sigma$ if
each matrix-valued polynomial $P_{j}$ is of degree $j$;
the matrix-valued polynomials $P_{j}$ satisfy the orthogonality relations:
$\int\limits_{[0,+\infty)}P_{j}(t)\,\sigma(dt)\,P_{k}^{*}(t)^{*}=\delta_{jk}C_{qj},$
for all $j,k\in\mathbb{N}$. Here, $\delta_{jk}$ is the Kronecker delta, and and $C_{qj}$ is a constant nonnegenative $q\times q$ matrix.
The following matrix polynomials seem to have been first introduced in [24].
Definition 2.8.
Let $(s_{j})_{j=0}^{2n}$ (resp. $(s_{j})_{j=0}^{2n+1}$) be a Stieltjes positive sequence. Furthermore, let $v_{j}$, $u_{1,j}$, $u_{2,j}$ and $R_{j}$ be as in (28), (32) and (25), respectively. Let
$P_{1,0}(z):=I_{q},\ \ Q_{1,0}(z):=0_{q},\ \ P_{2,0}(z):=I_{q},\ \ Q_{2,0}(z):=s_{0}.$ (38)
For $j\geq 1$, set
$P_{1,j}(z):=$ $(-Y_{1,j}^{*}H_{1,j-1}^{-1},\,I_{q})R_{j}(z)v_{j},$ (39)
$P_{2,j}(z):=$ $(-Y_{2,j}^{*}H_{2,j-1}^{-1},\,I_{q})R_{j}(z)v_{j},$ (40)
$Q_{1,j}(z):=$ $-(-Y_{1,j}^{*}H_{1,j-1}^{-1},\,I_{q})R_{j}(z)u_{1,j},$ (41)
$Q_{2,j}(z):=$ $-(-Y_{2,j}^{*}H_{2,j-1}^{-1},\,I_{q})R_{j}(z)u_{2,j}.$ (42)
The matrix polynomials $Q_{1,j}$ and $Q_{2,j}$ are called second-kind polynomials with respect to $P_{1,j}$ and $P_{2,j}$, respectively.
The matrix polynomials $P_{1,j}$ and $P_{2,j}$ are orthogonal with respect to a positive measure on $[0,+\infty)$. Simultaneously, the second-kind polynomials $Q_{1,j}$ and $Q_{2,j}$ can be derived from the corresponding polynomials $P_{1,j}$ and $P_{2,j}$, respectively.
Proposition 2.9.
a) The polynomials $P_{1,j}$ and $P_{2,j}$ are OMP with respect to $\sigma$ and $t\sigma$, respectively. More precisely (as seen in [17, Remark 6.3]),
$\int\limits_{[0,+\infty)}P_{k,j}(t)t^{k-1}\sigma(dt)P_{k,l}^{*}(t)=\left\{\begin{array}[]{ccc}0_{q}&:&j\neq l\\ \widehat{H}_{k,j}&:&j=l\end{array}\right.,\ \ k=1,2.$ (43)
b) The following identities hold (as seen in [7, Remark D.6] and [7, Remark E.4]):
$Q_{1,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{P_{1,j}(x)-P_{1,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n,$ (44)
$Q_{2,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{xP_{2,j}(x)-tP_{2,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n-1.$ (45)
By using the polynomials $P_{k,j}$ and $Q_{k,j}$ for $k=1,2$, we define the so-called resolvent matrix of the TSMM problem.
Definition 2.10.
[7, Theorem 4.7] For $k=0,1$, let $P_{k,j}$ and $Q_{k,j}$ be defined as in Definition 2.8. Define the $2q\times 2q$ matrix-valued function
$U^{(2m+k)}(z):=\left(\begin{array}[]{cc}Q_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)&-Q_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\\ -zP_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)&P_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\\ \end{array}\right),\quad k=0,1.$ (46)
For $k=0$ (resp. $k=1$), the matrix (46) is called the resolvent matrix of the TSMM problem for an odd number of moments $\left(s_{j}\right)_{j=0}^{2m}$ (resp. an even number of moments $\left(s_{j}\right)_{j=0}^{2m+1}$). In (46), the superscript of $U^{(2m+k)}$ for $k=0,1$ indicates the subscript of the moment $s_{2m+k}$ that appears in (46).
In [23], the identity
$U^{(2j)}(z)=$ $\left(\begin{array}[]{cc}I_{q}&0_{q}\\ -zM_{0}&I_{q}\end{array}\right)\left(\begin{array}[]{cc}I_{q}&L_{0}\\ 0&I_{q}\\ \end{array}\right)\ldots\left(\begin{array}[]{cc}I_{q}&L_{j-1}\\ 0_{q}&I_{q}\\ \end{array}\right)\left(\begin{array}[]{cc}I_{q}&0_{q}\\ -zM_{j}&I_{q}\end{array}\right)$ (55)
was verified and was used to express the solution of the TSMM problem through a finite matrix continued fraction. See the right-hand side of (60) and (61).
The complete set of solutions of the TSMM problem in the nondegenerate case is provided by the linear fractional transformation [17, Theorem 4.3]
$s^{(2m+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}(z):=\frac{Q_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)\,{{\bf p}}(z)-Q_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\,{{\bf q}}(z)}{-zP_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)\,{{\bf p}}(z)+P_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\,{{\bf q}}(z)},\,k=0,1,$ (56)
where ${\bf p}$ and ${\bf q}$ are certain meromorphic matrix-valued $q\times q$ functions; see [17, page 12]. The superscript of $s^{(2n+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}$ for $k=0,1$ indicates the subscript of the moment $s_{2m+k}$ appearing in (46), and the subscript of (56) emphasizes the dependency on the pair $\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}$ for the associated solution $s^{(2m+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}$ to the TSMM problem.
By choosing the pair $\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)$ equal to $\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)$ or $\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)$, we attain the so-called extremal solutions to the TSMM problem. In particular, we will consider the following extremal solutions of the TSMM problem:
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z):=-\frac{Q_{2,m}^{*}(\bar{z})}{z\,P_{2,m}^{*}(\bar{z})}$ (57)
and
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z):=-\frac{Q_{1,m}^{*}(\bar{z})}{P_{1,m}^{*}(\bar{z})}.$ (58)
By using [23] and an equivalent form of the resolvent matrix $U^{(2m)}$, it is proven that
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(x)<s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(x),\quad x\in(-\infty,0).$ (59)
In [7, Theorem 3.4], a representation of extremal solutions (58) and (57) in terms of matrix continued fractions was given. We reproduce this result in the following remark.
Remark 2.11.
Let $P_{k,j}$, $Q_{k,j}$, for $k=1,2$ be as in Definition 2.8 and for $M_{k}$, $L_{k}$ as in (35) through (37). The extremal solutions to the TSMM problem defined in $z\in\mathbb{C}\setminus[0,\infty)$ then take the following form:
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z)=-\frac{Q_{2,m}^{*}(\bar{z})}{zP_{2,m}^{*}(\bar{z})}=$ $\cfrac{I_{q}}{-zM_{0}+\cfrac{I_{q}}{L_{0}+\cfrac{I_{q}}{\ddots\,-zM_{m-1}+\cfrac{I_{q}}{L_{m-1}-z^{-1}M_{m}^{-1}}}}},$ (60)
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z)=-\frac{Q_{1,m}^{*}(\bar{z})}{P_{1,m}^{*}(\bar{z})}=$ $\cfrac{I_{q}}{-zM_{0}+\cfrac{I}{L_{0}+\cfrac{I_{q}}{\ddots+L_{m-2}+\cfrac{I_{q}}{-zM_{m-1}+L_{m-1}^{-1}}}}}.$ (61)
2.4 Relatively prime matrix polynomials
Recall representation (2) of an HTM polynomial ${\bf f}_{n}$. To explain that the pairs $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ (resp. $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$) are coprime, as in Theorem 3.5, we reproduce Definition 1.15.7 and Theorem 1.15.5 of [44] related to relatively right prime or right coprime $q\times q$ matrix polynomials.
Definition 2.12.
a) A $q\times q$ matrix $W(z)$ is called unimodular if $\det W(z)$ is equal to a nonzero constant quantity [45, page 370], independent of $z$.
b) Two $q\times q$ polynomial matrices $A(z)$ and $B(z)$ are called left (right) or row (column) equivalent if and only if one can be obtained from the other as a result of a finite number of elementary operations performed on its rows (columns). Specifically, $B(z)=L(z)A(z)$ (resp. $B(z)=A(z)P(z)$) where $L(z)$ (resp. $P(z)$) is the product of unimodular matrices corresponding to elementary operations on rows (columns). See [44, Definition 1.15.7].
Definition 2.13.
A matrix $L(z)\in\mathbb{C}^{q\times q}$ is called a left common divisor of $q\times q$ matrices $A(z)$ and $B(z)$ if and only if there exist matrices $A_{1}(z),B_{1}(z)\in\mathbb{C}^{q\times q}$ such that
$A(z)=L(z)A_{1}(z)\qquad\mbox{and}\qquad B(z)=L(z)B_{1}(z).$ (62)
A matrix $P(z)\in\mathbb{C}^{q\times q}$ is called a right common divisor of $q\times q$ matrices $A(z)$ and $B(z)$ if and only if there exist matrices $A_{2}(z),B_{2}(z)\in\mathbb{C}^{q\times q}$ such that
$A(z)=A_{2}(z)P(z)\qquad\mbox{and}\qquad B(z)=B_{1}(z)P(z).$ (63)
Definition 2.14.
Let $q\times q$ matrix-valued $A(z)$ and $B(z)$ are called relatively left prime (resp. right prime) if and only if only their left (resp. right) common divisors are unimodular matrices.
The following theorem appears in [44, Theorem 1.15.5], [61, Theorem 6.1] and [66, Theorem 2.5.16].
Theorem 2.15.
[…]
Matrix polynomials $A(z)$, $B(z)\in\mathbb{C}^{q\times q}$ are right coprime if and only if $\begin{pmatrix}A(z)\\ B(z)\end{pmatrix}$ and $\begin{pmatrix}I_{q}\\ 0_{q}\end{pmatrix}$ are left equivalent.
Related to the coprimeness of two matrix polynomials, we reproduce [45, Lemma 6.3-5], [3, page 585], [52, page 26].
Lemma 2.16.
Matrices $A(z)$ and $B(z)$ are right coprime if and only if polynomial matrices $X(z)$ and $Y(z)$ exist such that
$X(z)A(z)+Y(z)B(z)=I.$ (64)
Equality (64) is called Bézout’s identity.
Identity (64) also appears in [28, page 21] and [65, Corollary 4.1.4] in the context of control systems theory.
2.5 The Bezoutian and inertia for matrix polynomials
In this subsection, we reproduce some notions related to the inertia for a matrix polynomial [53, 54]. The inertia for scalar polynomials was considered in [51, page 462].
Let $L(\lambda)=L_{0}\lambda^{\ell}+L_{1}\lambda^{\ell-1}+\ldots+\lambda L_{\ell-1}+L_{\ell}$ be a $q\times q$ matrix polynomial with coefficients $L_{k}\in\mathbb{C}^{q\times q}$. If $L_{0}\neq 0_{q}$, then ${\rm deg}(L)=\ell$ is the degree of polynomial $L$. Furthermore, if $L_{0}=I_{q}$, then $L$ is called monic. Let $M_{1},L_{1},M$ and $L$ be polynomial matrices such that $M_{1}(\lambda)L_{1}(\lambda)-M(\lambda)L(\lambda)=0$. The Bezoutian associated with the quadruple $(M_{1},L_{1},M,L)$ is the block matrix
$B_{M_{1},M}(L,L_{1})=\left(\Gamma_{j,k}\right)_{j,k=1}^{m-1,\ell-1}.$ (65)
The block entries $\Gamma_{j,k}$ of the quadruple $(M_{1},L_{1},M,L)$ are determined by following equality:
$\sum_{j,k=0}^{m=1,\ell-1}\Gamma_{jk}\lambda^{j}\mu^{k}=\frac{1}{\lambda-\mu}\left(M_{1}(\lambda)L_{1}(\mu)-M(\lambda)L(\mu)\right).$ (66)
The quantity $\ell$ (resp. $m$) denotes the maximal degree of $L(\lambda)$, $L_{1}(\lambda)$ (resp. $M(\lambda)$, $M_{1}(\lambda)$). Assume that there exists $\lambda_{0}\in\mathbb{C}$ such that $\det L(\lambda_{0})\neq 0$: on this case $L$ is called regular.
Let $L$ be a regular polynomial of degree $\ell$. The triple
$\gamma(L)=\left(\gamma_{+}(L),\gamma_{-}(L),\gamma_{0}(L)\right)$ (67)
is called inertia with respect to the real axis, where $\gamma_{+}(L)$ (resp. $\gamma_{-}(L)$) is the number of zeros of $\det(L(\lambda))$ in the open upper (resp. open lower) half-plane, and $\gamma_{0}(L)$ is the number of zeros of $\det(L(\lambda))$ in the real axis. Now we reproduce [53, Corollary 2.2].
Corollary 2.17.
Let $L(\lambda)$ be a matrix polynomial with an invertible leading coefficient. The spectrum of $L(\lambda)$ lies in the upper half-plane if and only if there exists a regular matrix polynomial $L_{1}(\lambda)$ that satisfies
$L_{1}^{*}(\bar{\lambda})L_{1}(\lambda)=L^{*}(\bar{\lambda})L(\lambda)$ (68)
such that the matrix
${\displaystyle\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})}$ (69)
is positive definite.
To compute the matrix (69), we use the generalized Bezoutian matrix and form introduced by Anderson and Jury in [2], which is a $q\times q$ rational function $W(z)$. Suppose that $W(z)$ admits two matrix factorizations:
$W(z)=A^{-1}(z)B(z)=D(z)C^{-1}(z).$ (70)
The generalized Bezoutian form associated with the quadruple ${A,B,C,D}$ is defined as
$\Gamma(x,y)=\frac{1}{x-y}\left(A(x)D(y)-B(x)G(y)\right).$ (71)
Note that in this work the relation (68) and the matrix (69) are associated with the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$), while the decomposition (70) is connected to the identity
$s_{{\bf f}_{n}}(z)=s_{{\bf f}_{n}}^{*}(\bar{z}),$ (72)
where $s_{{\bf f}_{n}}(z)$ denotes the left-hand side of (10), (11). See Remark 3.1.
In the following remark, we reproduce an assertion from [2, Remark 2.2] concerning the integrability of the rational function $\Gamma$ in (70): the conditions under which $\Gamma$ is a polynomial in both $x$ and $y$.
Remark 2.18.
For $\Gamma(x,y)$ to be a polynomial in both $x$ and $y$, it is necessary and sufficient that
$A(x)D(x)=B(x)C(x).$ (73)
Our strategy for analyzing the Hurwitzness of the HTM polynomials is as follows: we explicitly compute the matrix (69) by using the representations given in (137) and (138) for the quadruple $({\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x))$. Rather than applying formula (87) directly, we employ Bezoutian forms associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$, corresponding to the cases $n=2m+1$ and $n=2m$, respectively. See Equalities (88) and (89).
3 Properties of the Hurwitz-type matrix polynomials
In this section, we consider properties of Hurwitz-type matrix polynomials.
Remark 3.1.
Let $s_{{\bf f}_{n}}(z)$ denote the left-hand side of (10), (11). By taking into account that ${\bf c}_{j}$ and ${\bf d}_{j}$ are positive definite, equality (72) is valid.
Theorem 3.2.
Let $n$ be a natural number such that $n\geq 2$. Furthermore, let ${\bf f}_{n}$ be a matrix-valued polynomial and ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) be the ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) parts of ${\bf f}_{n}$. The assertions below are valid.
  • (a)
    The HTM polynomial ${\bf f}_{2m}$ (resp. ${\bf f}_{2m+1}$) is associated with the finite sequence of Markov parameters or moments $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$), which are Hermitian matrices.
  • (b)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) be constructed via Hermitian Markov parameters of the polynomial ${\bf f}_{n}$. The matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) are positive definite matrices.
  • (c)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) be the Schur complements of Hermitian block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) , respectively. Thus, ${\bf f}_{n}$ is an HTM polynomial if and only if the Schur complements $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) are positive definite matrices. The matrices $\widehat{H}_{1,m-1}$ and $\widehat{H}_{2,m-1}$ (resp. $\widehat{H}_{1,m}$ and $\widehat{H}_{2,m-1}$) are constructed via the Markov parameters of the polynomial ${\bf f}_{n}$.
  • (d)
    Let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$. The matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the rational matrix-valued polynomial $s_{{\bf f}_{n}}(-z)$, as in (72), is an extremal solution to the TSMM problem, as in (58) (resp. (57)).
  • (e)
    Let $n=2m$ (resp. $n=2m+1$). The pair of sequences $(({\bf c}_{k})_{k=0}^{m},({\bf d}_{k})_{k=0}^{m-1})$ $(resp.(({\bf c}_{k})_{k=0}^{m},({\bf d}_{k})_{k=0}^{m}))$ as in Definition 1.1 A and A the pair of sequences $((M_{k})_{k=0}^{m},(L_{k})_{k=0}^{m-1})$ $(resp.((M_{k})_{k=0}^{m},(L_{k})_{k=0}^{m}))$ as in (35)-(37) coincide.
  • (f)
    For $n\geq 2$, the matrix polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if ${\bf f}_{n}$ has the following representation:
    ${\bf f}_{n}(z)=\left\{\begin{array}[]{cl}(-1)^{m}(P_{1,m}^{*}(-\bar{z}^{2})-z\,Q_{1,m}^{*}(-\bar{z}^{2})),&n=2m,\\ (-1)^{m}(Q_{2,m}^{*}(-\bar{z}^{2})+z\,P_{2,m}^{*}(-\bar{z}^{2})),&n=2m+1,\end{array}\right.$ (74)
    where $P_{k,m}$ and $Q_{k,m}$ for $k=1,2$ are polynomials defined in Definition 2.8. The matrix polynomials $P_{k,m}$ and $Q_{k,m}$ are constructed via the Markov parameters of the polynomial ${\bf f}_{n}$.
Proof.
Part (a) is proven in [8, Theorem 7.9]. Part (b) is proven in [8, Theorem 7.10]. Part (c) is established using part (b) and [36, Corollary 1], which demonstrate that the positive definiteness of the block matrix $H_{k,j}$ is equivalent to that of its Schur complement $\widehat{H}_{k,j}$. Part (d) follows from [7, Theorem 7.12] and the uniqueness of the extremal solutions (61) and (60). Part (e) is proven in [8, Theorem 7.12]. The proof of part (f) follows directly from (2), part (b) and the equalities
${\bf h}_{n}(z)=$ $(-1)^{m}P_{1,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m+1}Q_{1,m}^{*}(-\bar{z}),\qquad n=2m$ (75)
${\bf h}_{n}(z)=$ $(-1)^{m}Q_{2,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m}P_{2,m}^{*}(-\bar{z}),\qquad n=2m+1.$ (76)
Equalities (75) and (76) were proven in the proof of [7, Proposition 7.8]. ∎
The corresponding distribution functions of the Stieltjes transforms $s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z)$ and $s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)}}(z)$, associated with the HTM polynomials as given in (60) and (61), can be represented by monotonically non-decreasing, piecewise constant matrix function [4].
A $q\times q$ matrix $\widetilde{\sigma}(\cdot)$ defined on $[0,\infty)$ is called monotonically nondecreasing if $\widetilde{\sigma}(t)$ is self-adjoint on $[0,\infty)$ and $\widetilde{\sigma}(t_{2})-\widetilde{\sigma}(t_{1})\geq 0$ for $t_{1}\leq t_{2}$. For $n=2m+1$ (resp. $n=2m$), the points of discontinuity of the distribution function $\widetilde{\sigma}$ that correspond to the measure $\sigma$ are given by the roots of polynomial $\det(zP_{2,m}^{*}(-\bar{z}))$ (resp. $\det(P_{1,m}^{*}(-\bar{z}))$). On the other hand, polynomials (75) and (76), which are constructed using the moments $s_{j}$, can be obtained from an absolutely continuous distribution on $[0,\infty)$ that corresponds to a positive measure on $[0,\infty)$, as shown in the next example.
Example 3.3.
Let $\widetilde{\sigma}(x)=\begin{pmatrix}4-2e^{-x/2}&ie^{-x}\\ -ie^{-x}&2-e^{-x}\\ \end{pmatrix}$ be defined on $[0,\infty)$. We construct the moments $s_{0}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix}$, $s_{1}=\begin{pmatrix}4&-i\\ i&1\end{pmatrix}$, $s_{2}=\begin{pmatrix}16&-2i\\ 2i&2\end{pmatrix},$ $s_{3}=\begin{pmatrix}96&-6i\\ 6i&6\end{pmatrix}$, $s_{4}=\begin{pmatrix}768&-24i\\ 24i&24\end{pmatrix}$ of $\widetilde{\sigma}(x)$. The sequence $(s_{j})_{j=0}^{4}$ is a Stieltjes positive definite sequence. By using these moments and Definition 2.8, we derive $P_{2,2}$ and $Q_{2,2}$. From (74), we obtain the HMT polynomial
${\bf f}_{5}(z)=z^{5}I_{2}+A_{1}z^{4}+A_{2}z^{3}+A_{3}z^{2}+A_{4}z+A_{5},$ (77)
where $A_{1}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix},A_{2}=\begin{pmatrix}\frac{486}{37}&0\\ -\frac{264}{37}i&6\end{pmatrix},A_{3}=\begin{pmatrix}\frac{560}{37}&-5i\\ 5i&5\end{pmatrix},$
$A_{4}=\begin{pmatrix}\frac{1158}{37}&0\\ -\frac{936}{37}i&6\\ \end{pmatrix}$ and $A_{5}=\begin{pmatrix}\frac{292}{37}&-2i\\ 2i&2\\ \end{pmatrix}.$ These coefficients $A_{j}$ clearly satisfy Condiction C (14). Therefore, by Theorem 5.3, the polynomial ${\bf f}_{5}$ is a Hurwitz polynomial.
3.1 Coprimeness of pairs $({\bf h}_{n}(z),g_{n}(z))$ and $({\bf h}_{n}(z),zg_{n}(z))$
We paraphrase [7, Proposition 5.1 (a)] to apply it in Theorem 3.5.
Proposition 3.4.
Let $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) be Stieltjes positive sequences. Furthermore, let $P_{k,j}$ and $Q_{k,j}$ be as in Definition 2.8. Moreover, let $\widehat{H}_{1,j}$ be as in (33). Thus, the following equalities hold:
$P_{1,j}(z)Q_{2,j}^{*}(\bar{z})-zQ_{1,j}(z)P_{2,j}^{*}(\bar{z})=$ $\widehat{H}_{1,j},$ (78)
$P_{k,j}(z)Q_{2,j}^{*}(\bar{z})-Q_{k,j}(z)P_{k,j}^{*}(\bar{z})=$ $0_{q}.$ (79)
In the following result, we consider the coprimeness of the ${\bf h}_{n}$ and ${\bf g}_{n}$ parts of a given polynomial ${\bf f}_{n}$. To prove this, we utilize Bézout’s identity (64), which involves orthogonal matrix polynomials $P_{k,j}$ on $[0,+\infty)$ and their second-kind polynomials $Q_{k,j}$.
Theorem 3.5.
(a) Let ${\bf f}_{2m}(z)$ be an HTM polynomial. Thus, its corresponding ${\bf h}_{2m}(z)$ and ${\bf g}_{2m}(z)$ are right coprime polynomials.
(b) Let ${\bf f}_{2m+1}(z)$ be an HTM polynomial. Thus, its corresponding ${\bf h}_{2m+1}(z)$ and $z{\bf g}_{2m+1}(z)$ are right coprime polynomials.
Proof.
From Theorem 3.2(a), we deduce that the polynomial ${\bf f}_{2m+1}$ (resp. ${\bf f}_{2m}$) is associated with the $(s_{j})_{j=0}^{2m}$ (resp $(s_{j})_{j=0}^{2m-1}$) Markov parameters. For the polynomial ${\bf f}_{2m}$, we apply Proposition 2.5 and find a parameter $s_{2m}$ that satisfies Inequality (21). We exploit the fact that $(s_{j})_{j=0}^{2m-1}$ is a Hankel nonnegative definite extendable sequence and that $s_{2m}$ represents the indicated extension. See Definition 2.3. Consequently, polynomials $P_{1,m}(z)$ and $Q_{2,m}^{*}(\bar{z})$ are well defined for ${\bf f}_{2m}$ and ${\bf f}_{2m+1}$. Let $n=2m+1$.
By using direct calculations and (76), we rewrite (78) for $j=m$ as
$X_{1}(z){\bf h}_{2m+1}(z)+Y_{1}(z)(z\,{\bf g}_{2m+1}(z))=I,$ (80)
where $X_{1}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}P_{1,m}(-z)$ and $Y_{1}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}Q_{1,m}(-z)$. Recall that the Schur complement $\widehat{H}_{1,m}$ is invertible because $H_{1,m}$ is positive definite.
For $n=2m$, by taking the adjoint of (78) and by replacing $z$ with $\bar{z}$, we obtain
$X_{2}(z){\bf h}_{2m}(z)+Y_{2}(z){\bf g}_{2m}(z)=I,$ (81)
where $X_{2}(z):=(-1)^{m}\widehat{H}_{1,m}^{-1}Q_{2,m}(-z)$ and $Y_{2}(z):=(-1)^{m}z\widehat{H}_{1,m}^{-1}P_{2,m}(-z)$.
From (80) (resp. (81) ) and Lemma 2.16, we observe that the polynomials ${\bf h}_{2m+1}(z)$ and $z\,{\bf g}_{2m+1}(z)$ (resp. ${\bf h}_{2m}(z)$ and ${\bf g}_{2m}(z)$) are right coprime. ∎
Now we present an alternative proof of Theorem 3.5. Although different regarding the inverse of the matrix $U^{(2m)}$ as in (46), a similar proof was provided in [12].
Alternative proof of Theorem 3.5. The inverse of the resolvent matrix (46) [7] is the following:
$U^{{(2m)}^{-1}}(z)=\begin{pmatrix}P_{1,m}^{-1}(0)P_{1,m}(z)&P_{1,m}^{-1}(0)Q_{1,m}(z)\\ zQ_{2,m}^{-1}(0)P_{2,m}(z)&Q_{2,m}^{-1}(0)Q_{2,m}(z)\end{pmatrix},\quad z\in\mathbb{C}.$ (82)
To prove (a), we introduce the matrix
$X_{3}(z):={\rm diag}(\widehat{H}_{1,m}^{-1}P_{1,m}(0),I_{q})\,U^{{(2m)}^{-1}}(z)\,{\rm diag}((-1)^{m}I_{q},(-1)^{m+1}I_{q}).$ (83)
With (78), (79) and (76), it can be readily verified that
$X_{3}(z)\begin{pmatrix}{\bf h}_{2m+1}(z)\\ z\,{\bf g}_{2m+1}(z)\end{pmatrix}=\begin{pmatrix}I_{q}\\ 0\end{pmatrix}.$ (84)
By (55) and Definition 2.12 a), the $2q\times 2q$ matrix-valued polynomial $X_{3}(z)$ is unimodular. The fact that the pair $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$ is left coprime follows from Theorem 2.15, and part (a) is proved.
Now we prove part (b). Let $K_{1}:=\begin{pmatrix}0&\widehat{H}_{1,m}^{-1}Q_{2,m}(0)\\ I&0\end{pmatrix}$. Furthermore, let
$X_{4}(z):=K_{1}U^{{(2m)}^{-1}}(z)\,{\rm diag}((-1)^{m}I,(-1)^{m}I).$ (85)
By incorporating (78), (79) and (75), it can be readily verified that
$X_{4}(z)\begin{pmatrix}{\bf h}_{2m}(z)\\ {\bf g}_{2m}(z)\end{pmatrix}=\begin{pmatrix}I_{q}\\ 0_{q}\end{pmatrix}.$ (86)
By offering similar arguments as in the proof of part (a), it follows that the pair $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ is left coprime. Thus, part (b) is proven.
4 Bezoutian and HTM polynomials
In this section, under the commutative-like Condition C (14), we derive an explicit form of the Bezoutian for the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$). Instead of working directly with the Bezoutian form
${\mathscr{F}}_{n}(x,y):=$ $\frac{{\bf f}_{n}^{*}(\bar{x}){\bf f}_{n}(-y)-{\bf f}_{n}^{*}(-\bar{x}){\bf f}_{n}(y)}{x-y},$ (87)
we use the Equality (2) and calculate the Bezoutian forms
${\mathscr{G}}_{n}^{(1)}(x,y):=$ $\frac{1}{x^{2}-y^{2}}\left(x^{2}{\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-y^{2}{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})\right)$ (88)
and
${\mathscr{G}}_{n}^{(2)}(x,y):=$ $\frac{xy}{x^{2}-y^{2}}\left({\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})\right)$ (89)
associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$. Recall that a similar decomposition of the Bezoutian for scalar polynomials was previously studied; see [48, page 293] and [1, page 257]. Following [2] to compute the Bezoutian forms ${\mathscr{G}}_{n}^{(1)}(x,y)$ and ${\mathscr{G}}_{n}^{(2)}(x,y)$ we employ both representations (72) of the rational matrices (10) and (11).
It readily follows that
${\mathscr{F}}_{n}(x,y)=2\left({\mathscr{G}}_{n}^{(1)}(x,y)+{\mathscr{G}}_{n}^{(2)}(x,y)\right).$ (90)
In the next lemma, we demonstrate that the matrix functions ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are matrix polynomials in $x$ and $y$. See [2, Remark 1].
Lemma 4.1.
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as defined in (87), (88) and (89), respectively. The functions ${\mathscr{G}}_{n}^{(1)}$, ${\mathscr{G}}_{n}^{(2)}$ and ${\mathscr{F}}_{n}$ are polynomials on $x$ and $y$.
Proof.
Let $n=2m$, and let
$\widetilde{V}_{[0,2m]}(x):=\left(\bar{x}^{2m}I_{q},\bar{x}^{2m-2}I_{q},\ldots,\bar{x}^{2}I_{q},I_{q}\right)^{\mathsf{*}}.$ (91)
For $i,j=0,2,4,\ldots,2m$, let
$\alpha_{ij}:=-A_{i}^{*}A_{j+1}+A_{i+1}^{*}A_{j},$ (92)
where $A_{j}$ is a coefficient of ${\bf f}_{n}$ and we set $A_{2m+1}=0$. Additionally, let
${\mathscr{A}}_{2m}:=\begin{pmatrix}0&\alpha_{0,2}&\alpha_{0,4}&\ldots&\alpha_{0,2m-2}&\alpha_{0,2m}\\ -\alpha_{0,2}&0&\alpha_{2,4}&\ldots&\alpha_{2,2m-2}&\alpha_{2,2m}\\ -\alpha_{0,4}&-\alpha_{2,4}&0&\ldots&\alpha_{4,2m-2}&\alpha_{4,2m}\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ -\alpha_{0,2m-2}&-\alpha_{2,2m-2}&-\alpha_{4,2m-2}&\ldots&0&\alpha_{2m-2,2m}\\ -\alpha_{0,2m}&-\alpha_{2,2m}&-\alpha_{4,2m}&\ldots&-\alpha_{2m-2,2m}&0\end{pmatrix}$ (93)
By using (5) , (9), (91) , (92), (93), as well as Condition C (14), we have
$x^{2}{\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-y^{2}{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})=\widetilde{V}_{[0,2m]}^{*}(\bar{x}){\mathscr{A}}_{2m}\widetilde{V}_{[0,2m]}(y)$
$=\alpha_{0,2m}(x^{2m}-y^{2m})+\alpha_{0,2m-2}x^{2}y^{2}(x^{2m-2}-y^{2m-2})+\ldots+\alpha_{2m-2,2m}(x^{2}-y^{2}).$
It follows that ${\mathscr{F}}_{n}$ is a polynomial in $x$ and $y$. From the last equality and right-hand side of (88), we see that ${\mathscr{G}}_{2m}^{(1)}(x,y)$ is polynomial on $x$ and $y$.
Similarly, it can be shown that ${\mathscr{G}}_{2m}^{(2)}$, ${\mathscr{G}}_{2m+1}^{(1)}$ and ${\mathscr{G}}_{2m+1}^{(2)}$ are polynomials in $x$ and $y$. From (90), it follows that ${\mathscr{F}}_{n}$ is a polynomial in $x$ and $y$. ∎
For $k=0,1$, let
$H^{(k)}_{\infty}:=$ $\begin{pmatrix}s_{k}&-s_{k+1}&s_{k+2}&\ldots\\ -s_{k+1}&s_{k+2}&\iddots\\ s_{k+2}&\iddots\end{pmatrix}$ (94)
be an infinite Hankel matrix.
A similar untruncated block Hankel matrix, denoted by $\bf{H}$, appears in the proof of [2, Lemma 2.3].
Lemma 4.2.
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as in (87), (88) and (89), respectively. Thus, for $n=2m$, we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-2}I_{q},y^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}),$ (95)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(0)}_{\infty}\left(y^{-1}I_{q},y^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}).$ (96)
For $n=2m+1$, we have
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ ${\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(I_{q},x^{-2}I_{q},\ldots\right)H^{(0)}_{\infty}\left(I_{q},y^{-2}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(y^{2}),$ (97)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-{\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-1}I_{q},y^{-3}I_{q},\ldots\right)^{\mathsf{T}}$
$\cdot{\bf g}_{2m+1}(y^{2}),$ (98)
where the superscript $\mathsf{T}$ signifies the transpose of a matrix.
Proof.
We prove (95). With (16), we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ $\frac{1}{x^{2}-y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{2}{\bf h}_{2m}^{-1^{*}}(\bar{x}^{2}){\bf g}_{2m}^{*}(\bar{x}^{2})-y^{2}{\bf g}_{2m}(y^{2}){\bf h}_{2m}^{-1}(y^{2})\right)$
$\cdot{\bf h}_{2m}(y^{2})$
$=$ $\frac{1}{x^{2}-y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(s_{1}(y^{-2}-x^{-2})-s_{2}(y^{-4}-x^{-4})+s_{3}(y^{-6}-x^{-6})\right.$
$\left.-\ldots\right){\bf h}_{2m}(y^{2})$
$=$ $\frac{1}{x^{2}y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(s_{1}-s_{2}(x^{-2}+y^{-2})+s_{3}(x^{-4}+x^{-2}y^{-2}+y^{-4})\right.$
$\left.-\ldots\right){\bf h}_{2m}(y^{2})$
$=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(y^{-2}I_{q},y^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(y^{2}).$
In the last equality, we substituted (94) for $k=1$. Similarly, to prove (96), we use (16) and (94) for $k=0$. Equalities (97) and (98) are proven by applying (17) and (94) for $k=0$ and $k=1$, respectively. ∎
Let
${\bf h}_{2m}^{(11)}(\xi)+{\bf h}_{2m}^{(12)}(\xi)=$ $\left(\xi^{-2}I_{q},\xi^{-4}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(\xi^{2}),$ (99)
${\bf h}_{2m}^{(21)}(\xi)+{\bf h}_{2m}^{(22)}(\xi)=$ $\left(\xi^{-1}I_{q},\xi^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf h}_{2m}(\xi^{2}),$ (100)
${\bf g}_{2m+1}^{(11)}(\xi)+{\bf g}_{2m+1}^{(12)}(\xi)=$ $\left(I_{q},\xi^{-2}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(\xi^{2}),$ (101)
${\bf g}_{2m+1}^{(21)}(\xi)+{\bf g}_{2m+1}^{(22)}(\xi)=$ $\left(\xi^{-1}I_{q},\xi^{-3}I_{q},\ldots\right)^{\mathsf{T}}{\bf g}_{2m+1}(\xi^{2}),$ (102)
where ${\bf h}_{2m}^{(k1)}(\xi)$ (resp. ${\bf g}_{2m+1}^{(k1)}(\xi)$) for $k=1,2$ are polynomials in $\xi$, and ${\bf h}_{2m}^{(k2)}(\xi)$ (resp. ${\bf g}_{2m+1}^{(k2)}(\xi)$) for $k=1,2$ are polynomials in $\xi^{-1}$.
For a natural number $m$, let
$\widetilde{J}_{m}:={\rm diag}\,\left(I_{q},(-1)I_{q},(-1)^{2}I_{q}\ldots,(-1)^{m}I_{q}\right)$ (103)
be a $(m+1)q\times(m+1)q$. The following remark readily follows from direct calculation.
Remark 4.3.
Let ${\bf h}_{2m}^{(11)}(\xi)$, ${\bf h}_{2m}^{(21)}(\xi)$, ${\bf g}_{2m+1}^{(11)}(\xi)$ and ${\bf g}_{2m+1}^{(21)}(\xi)$ be as defined in (99), (100), (101) and (102), respectively. Furthermore, let $H_{1,j}$, $H_{2,j}$ and $J_{m}^{(k)}$ be as defined in (18) and (103), respectively. Thus, the following equalities hold
${\bf h}_{2m}^{(11)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf h}_{2m}^{(11)}(y)=$ $\widehat{\bf h}_{2m}^{(11)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(11)}(y),$ (104)
${\bf h}_{2m}^{(21)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf h}_{2m}^{(21)}(y)=$ $\widehat{\bf h}_{2m}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(21)}(y),$ (105)
${\bf g}_{2m+1}^{(11)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf g}_{2m+1}^{(11)}(y)=$ $\widehat{\bf g}_{2m+1}^{(11)^{*}}(\bar{x})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}\,\widehat{\bf g}_{2m+1}^{(11)}(y),$ (106)
${\bf g}_{2m+1}^{(21)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf g}_{2m+1}^{(21)}(y)=$ $\widehat{\bf g}_{2m+1}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf g}_{2m+1}^{(21)}(y),$ (107)
where
$\widehat{\bf h}_{2m}^{(11)},\,\widehat{\bf h}_{2m}^{(21)},\,\widehat{\bf g}_{2m+1}^{(11)}\quad\mbox{and}\quad\widehat{\bf g}_{2m+1}^{(11)}$ (108)
represents finite-dimensional restrictions of the matrices ${\bf h}_{2m}^{(11)}$, ${\bf h}_{2m}^{(21)}$, ${\bf g}_{2m+1}^{(11)}$ and ${\bf g}_{2m+1}^{(11)}$. In particular, the polynomial part of
$\widehat{\bf h}_{2m}^{(11)}(y)=\left(y^{-2}{\bf h}_{2m}(y^{2}),y^{-4}{\bf h}_{2m}(y^{2}),\ldots,y^{-2m}{\bf h}_{2m}(y^{2})\right)^{\mathsf{T}}$ (109)
is the restriction of the polynomial part of the matrix
${\bf h}_{2m}^{(11)}(y)=\left(y^{-2}{\bf h}_{2m}(y^{2}),y^{-4}{\bf h}_{2m}(y^{2}),\ldots,y^{-2m}{\bf h}_{2m}(y^{2}),0_{q},0_{q}\ldots,\right)^{\mathsf{T}}.$
Equalities (104)–(107) correspond to those in the proof of [2, Lemma 2.3], where, instead of the matrices $H_{1,j}$ and $H_{2,j}$, the truncated block Hankel matrix ${\bf H}_{nm}$ is used. Moreover, in place of the matrix $H_{\infty}^{(1)}$, the infinite Hankel matrix $H$ is employed.
For a matrix polynomial $Q(x)=Q_{0}x^{\ell}+Q_{1}x^{\ell-1}+\ldots+Q_{\ell-1}x+Q_{\ell}$, the Hankel block matrix
$S(Q):=\begin{pmatrix}Q_{\ell-1}&Q_{\ell-2}&\ldots&Q_{1}&Q_{0}\\ Q_{\ell-2}&Q_{\ell-3}&\ldots&Q_{0}&0_{q}\\ \vdots&\iddots&\iddots&\iddots&\vdots\\ Q_{0}&0_{q}&\ldots&0_{q}&0_{q}\end{pmatrix}$ (110)
is called a symmetrizer of $Q(x)$. In [51, page 455], the concept of symmetrizer is introduced for scalar polynomials. Additionally, let
$F_{0}^{2k}(x):=$ $\left(I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2k}I_{q}\right)^{*},$ (111)
$F_{1}^{2k-1}(x):=$ $\left(\bar{x}I_{q},\bar{x}^{3}I_{q},\ldots,\bar{x}^{2k-1}I_{q}\right)^{*}.$ (112)
In the following theorem, added by symmetrizers of corresponding matrices and the block Hankel matrices $H_{k,j}$, we obtain a factorized expression of the Bezoutian forms ${\mathscr{G}}_{2m}^{(1)}$, ${\mathscr{G}}_{2m}^{(2)}$, ${\mathscr{G}}_{2m+1}^{(1)}$ and ${\mathscr{G}}_{2m+1}^{(2)}$.
Theorem 4.4.
Let Condition C, given by (14), be satisfied. Furthermore, let ${\mathscr{G}}_{2m}^{(k)}$, ${\mathscr{G}}_{2m+1}^{(k)}$ for $k=1,2$ be as defined in (88) and (89). Additionally, let ${\bf h}_{n}$, ${\bf g}_{n}$, $S$, $\widetilde{J}_{m}$, $H_{1,j}$, $H_{2,j}$, $F_{0}^{2k}(x)$ and $F_{1}^{2k-1}(x)$ be as in (5), (9), (110), (103), (18), (111) and (112), respectively.
Therefore, for $n=2m$, the following equality holds:
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ $F_{0}^{(2m-2)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{0}^{(2m-2)}(y),$ (113)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{1}^{(2m-1)}(y).$ (114)
For $n=2m+1$, we have
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ $F_{0}^{(2m)^{*}}(\bar{x})S({\bf g}_{2m+1}^{*})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}S({\bf g}_{2m+1})F_{0}^{(2m)}(y).$ (115)
Moreover,
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S(\widetilde{\bf g}_{2m+1}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S(\widetilde{\bf g}_{2m+1})F_{1}^{(2m-1)}(y),$ (116)
where
$\widetilde{\bf g}_{2m+1}(z):=A_{0}z^{m-1}+A_{2}z^{m-2}+\ldots+A_{2m-2}.$ (117)
Proof.
We prove (113). From (95), (99) and (104), we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=\widehat{\bf h}_{2m}^{(11)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(11)}(y).$ (118)
We examine $\widehat{\bf h}_{2m}^{(11)}(y)$. By using (109) and (110), we have
$\widehat{\bf h}_{2m}^{(11)}(y)=$ $\begin{pmatrix}A_{2m-2}+\ldots+A_{2}y^{2m-4}+I_{q}y^{2m-2}\\ A_{2m-4}+\ldots+A_{2}y^{2m-6}+I_{q}y^{2m-4}\\ \ldots\\ I_{q}\end{pmatrix}$
$=$ $\begin{pmatrix}A_{2m-2}&\ldots&A_{2}&I_{q}\\ A_{2m-4}&\ldots&I_{q}&\\ \ldots\\ I_{q}\end{pmatrix}\begin{pmatrix}I_{q}\\ y^{2}I_{q}\\ \ldots\\ y^{2m-2}I_{q}\end{pmatrix}$
$=$ $S({\bf h}_{2m})\left(I_{q},\bar{y}^{2}I_{q},\ldots,\bar{y}^{2m-2}I_{q}\right)^{*}.$ (119)
By substituting (119) for (118), we obtain (113). Similarly, by using
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $\widehat{\bf h}_{2m}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(21)}(y),$ (120)
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ $\widehat{\bf g}_{2m+1}^{(11)^{*}}(\bar{x})\widetilde{J}_{m}H_{1,m-1}\widetilde{J}_{m}\,\widehat{\bf g}_{2m+1}^{(11)}(y),$ (121)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $\widehat{\bf g}_{2m+1}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf g}_{2m+1}^{(21)}(y),$ (122)
one can prove (114), (115) and (116). Equalities (120), (121) and (122) readily follow from (105), (106), (107) and (100), as well as from (101) and (102). Additionally, one also uses (96), (97), (98), (111) and (112). ∎
In the following corollary, we represent ${\mathscr{F}}_{n}$ as a matrix quadratic form regarding the block Hankel matrices $H_{1,j}$ and $H_{2,j}$, as well as the symmetrizer matrices $S({\bf h}_{2m})$, $S({\bf g}_{2m+1})$ and $S(\widetilde{\bf g}_{2m+1})$.
Corollary 4.5.
Assume the same conditions and notations as in Theorem 4.4. Let $\widetilde{J}_{m}$ be as in (103). Moreover, let
${\mathscr{J}}_{r}^{(\ell)}:=\begin{pmatrix}\widetilde{J}_{\ell}&0\\ 0&\widetilde{J}_{r}\end{pmatrix}$ (123)
for all $r,\,\ell>1$. Furthermore, for $k=0,1$ let
${\mathscr{W}}_{2m}^{(k)}:=$ $\begin{pmatrix}S({\bf h}_{2m}^{*})&0\\ 0&S({\bf h}_{2m}^{*})\end{pmatrix}{\mathscr{J}}_{m-1}^{(m-1)}\begin{pmatrix}H_{2,m-1}&0\\ 0&(-1)^{k}H_{1,m-1}\end{pmatrix}{\mathscr{J}}_{m-1}^{(m-1)}$
$\cdot\begin{pmatrix}S({\bf h}_{2m})&0\\ 0&S({\bf h}_{2m})\end{pmatrix},$ (124)
${\mathscr{W}}_{2m+1}^{(k)}:=$ $\begin{pmatrix}S({\bf g}_{2m+1}^{*})&0\\ 0&S(\widetilde{\bf g}_{2m+1}^{*})\end{pmatrix}{\mathscr{J}}_{m-1}^{(m)}\begin{pmatrix}H_{1,m}&0\\ 0&(-1)^{k}H_{2,m-1}\end{pmatrix}{\mathscr{J}}_{m-1}^{(m)}$
$\cdot\begin{pmatrix}S({\bf g}_{2m+1})&0\\ 0&S(\widetilde{\bf g}_{2m+1})\end{pmatrix}.$ (125)
With ${\mathscr{F}}_{n}$ be as in (87), the following equalities hold:
${\mathscr{F}}_{2m}(x,y)=$ $2\begin{pmatrix}F_{0}^{(2m-2)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{W}}_{2m}^{(1)}\begin{pmatrix}F_{0}^{(2m-2)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix},$ (126)
and
${\mathscr{F}}_{2m+1}(x,y)=$ $2\begin{pmatrix}F_{0}^{(2m)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{W}}_{2m+1}^{(1)}\begin{pmatrix}F_{0}^{(2m)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}.$ (127)
Proof.
Equality (126) directly follows (90), (113), (114) and (124). In a similar manner, Equality (127) can readily be derived from (90), (115), (116) and (125). ∎
5 Hurwitness of Hurwitz-type matrix polynomials
In this section, we verify that an HTM polynomial ${\bf f}_{n}$, defined in Definition 1.1 and satisfying Condition C given by (14), is indeed a Hurwitz matrix polynomial. In [67], the Hurwitzness of HTM polynomials was considered. Unfortunately, the proof of [67, Theorem 3.5] regarding the inertia of the matrix polynomial ${\bf f}_{n}$ for $n=2m$ is incorrect, and the proof for $n=2m+1$ is absent. See Remark 5.8. Recall that the inertia of matrix polynomials, denoted by $\gamma({\bf f}_{n})$, provides the number of eigenvalues of ${\bf f}_{n}$ with positive, negative and real zero parts, respectively; see [53, page 410]. We rewrite Corollary 4.5 for the pair $(ix,iy)$ instead of $(x,y)$.
Corollary 5.1.
Assume the same conditions and notations as in Theorem 4.4. Moreover, let ${\mathscr{F}}_{n}$, ${\mathscr{J}}_{k}^{\ell}$, ${\mathscr{W}}_{2m}^{(0)}$ and ${\mathscr{W}}_{2m+1}^{(0)}$ be as in (87), (123), (124) and (125), respectively. Thus, the following equalities hold:
${\mathscr{F}}_{2m}(ix,iy)=$ $2\begin{pmatrix}F_{0}^{(2m-2)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}\begin{pmatrix}F_{0}^{(2m-2)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}$ (128)
and
${\mathscr{F}}_{2m+1}(ix,iy)=$ $2\begin{pmatrix}F_{0}^{(2m)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}\begin{pmatrix}F_{0}^{(2m)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}.$ (129)
Proof.
Equalities (128) and (129) readily follow from (126) and (127), respectively. ∎
Let
$V_{[0,2m-1]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-2}I_{q},\bar{x}^{2m-1}I_{q}\right)^{\mathsf{*}},$ (130)
$V_{[0,2m]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-1}I_{q},\bar{x}^{2m}I_{q}\right)^{\mathsf{*}}.$ (131)
Furthermore, let
$\begin{pmatrix}F_{0}^{(2m-2)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}V_{[0,2m-1]}(x)$ (132)
and
$\begin{pmatrix}F_{0}^{(2m)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(x),$ (133)
where ${\mathscr{T}}_{2m}$ (resp. ${\mathscr{T}}_{2m+1}$) are elementary block matrices. The following remark readily follows.
Remark 5.2.
Equalities (128) and (129) can be written as follows:
${\mathscr{F}}_{2m}(ix,iy)=$ $2V_{[0,2m-1]}^{*}(\bar{x}){\mathscr{T}}_{2m}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}V_{[0,2m-1]}(y),$ (134)
${\mathscr{F}}_{2m+1}(ix,iy)=$ $2V_{[0,2m]}^{*}(\bar{x}){\mathscr{T}}_{2m+1}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(y).$ (135)
We now state the main result of the present work.
Theorem 5.3.
Each HTM polynomial satisfying Condition C (14) is a Hurwitz matrix polynomial.
Proof.
We use Corollary 2.17 with
$L(\lambda)={\bf f}_{n}(i\lambda)\quad\mbox{and}\quad L_{1}(\lambda)={\bf f}_{n}(-i\lambda)$ (136)
and Equalities (65) , (66), (87), (88), (89) and (90). For $n=2m$, by using (134), we have
${\displaystyle\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})}=2{\mathscr{T}}_{2m}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}.$ (137)
By employing (135) for $n=2m+1$, we have
${\displaystyle\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})}=2{\mathscr{T}}_{2m+1}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}.$ (138)
The right-hand sides of (137) and (138) are positive definite matrices because
$\begin{pmatrix}H_{2,m-1}&0\\ 0&H_{1,m-1}\end{pmatrix}\quad\mbox{and}\quad\begin{pmatrix}H_{1,m}&0\\ 0&H_{2,m-1}\end{pmatrix}$
are positive definite, and both ${\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}$ and ${\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}$ are invertible matrices
From Corollary 2.17, the spectrum of HTM polynomials ${\bf f}_{2m}(i\lambda)$ and ${\bf f}_{2m+1}(i\lambda)$ belongs to the upper half-plane of ${\mathbb{C}}$. Consequently, the spectrum of ${\bf f}_{2m}(\lambda)$ and ${\bf f}_{2m+1}(\lambda)$ lies in the left half-plane of ${\mathbb{C}}$. Thus, polynomials ${\bf f}_{2m}(\lambda)$ and ${\bf f}_{2m+1}(\lambda)$ are Hurwitz matrix polynomials. ∎
Remark 5.4.
A $q\times q$ HTM polynomial ${\bf f}_{2}$ of degree $2$ is a Hurwitz matrix polynomial, since ${\bf f}_{2}$ admits the presentation
${\bf f}_{2}(z)=I_{q}z^{2}+s_{0}z+s_{0}^{-1}s_{1}$
and it readily satisfies Condition C (14).
Remark 5.5.
By employing (67), note that $\gamma_{-}({\bf f}_{n}(i\lambda))=\gamma_{0}({\bf f}_{n}(i\lambda))=0$ and $\gamma({\bf f}_{n}(i\lambda))=\gamma_{+}({\bf f}_{n}(i\lambda))$.
Proposition 5.6.
Let ${\bf f}_{n}$ be an HTM polynomial. Moreover, let ${\bf h}_{n}$ and ${\bf g}_{n}$ be as in (2) , (5) and (9), respectively. Furthermore, let $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ and $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ be Bezoutians associated with the quadruple $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$, respectively. Thus,
a) the matrix $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ is a positive definite, and
b) the matrix $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ is a negative definite.
Proof.
Let $\widetilde{J}_{m}$, $\widetilde{\bf g}_{2m+1}$, $S({\bf h}_{2m})$, $S({\bf g}_{2m+1})$ and $S(\widetilde{\bf g}_{2m+1})$ be as in (103), (117) and (110). For $n=2m$, by using (88), (89), (113) and (114), we have
$B_{x{\bf g}_{2m}^{*},{\bf h}_{2m}^{*}}(x{\bf g}_{2m},{\bf h}_{2m})=$ $S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m}),$
$B_{{\bf g}_{2m}^{*},{\bf h}_{2m}^{*}}({\bf g}_{2m},{\bf h}_{2m})=$ $-S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m}).$
By (88), (89), (115) and (116), for $n=2m+1$, we obtain
$B_{x{\bf g}_{2m+1}^{*},{\bf h}_{2m+1}^{*}}(x{\bf g}_{2m+1},{\bf h}_{2m+1})=$ $S({\bf g}_{2m+1}^{*})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}S({\bf g}_{2m+1})$
$B_{{\bf g}_{2m+1}^{*},{\bf h}_{2m+1}^{*}}({\bf g}_{2m+1},{\bf h}_{2m+1})=$ $-S(\widetilde{\bf g}_{2m+1}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S(\widetilde{\bf g}_{2m+1}).$
From Theorem 3.2 (b), see that assertion a) and b) are valid. ∎
Remark 5.7.
In [51, Theorem 2], it is proven that a scalar polynomial $f_{n}$ is Hurwitz if and only if $B_{xg_{n}^{*},h_{n}^{*}}(xg_{n},h_{n})$ is negative definite and $B_{g_{n}^{*},h_{n}^{*}}(g_{n},h_{n})$ is positive definite. Note that the Markov parameters related to [51, Theorem 2] have the opposite sign compared with those associated with parts (a) and (b) of Proposition 5.6.
Remark 5.8.
Let $L_{1}$ and $L$ be as in (136). In the proof of [67, Theorem 3.5] regarding the inertia representation of matrix polynomials ${\bf f}_{n}$, the [67, Equality (3.6)]
$B_{L_{1}^{*},L^{*}}(L,L_{1})=2iB_{{\bf h}_{2m}^{*}(-\bar{x}^{2}),x{\bf g}_{2m}^{*}(-\bar{x}^{2})}({\bf h}_{2m}(-x^{2}),-x{\bf g}_{2m}(-x^{2}))$ (139)
is employed for $n=2m$. Furthermore, from (137) and (124), we see that (139) is not correct. The proof of [67, Theorem 3.5] for $n=2m+1$ is not given.
In Theorem 4.10 of [67], a “stability criterion via Markov parameters” for matrix polynomials ${\bf f}_{n}$ in even and odd cases is suggested. We have the following remarks regarding this.
Remark 5.9.
a) The family of matrix polynomials considered in [67, Theorem 4.10] is related to the HTM polynomials. Other matrix polynomials, for example, the Hurwitz polynomial (143), are not covered by [67, Theorem 4.10].
b) The proof of [67, Theorem 4.10] uses a decomposition of a certain Bezoutian form as a polynomial in $z$ and $u$, without verifying whether Equality (73) is satisfied.
c) The proof of [67, Theorem 4.10] separately analyzes decompositions of the Bezoutian forms related to the odd part $F_{o}$ of the matrix polynomial $F$, and the even part $F_{e}$ of $F$, without indicating the relationship between the Bezoutian form of $F$ and those of $F_{o}$ and $F_{e}$.
d) The proof of [67, Theorem 4.10] is performed out for the case $n=2m+1$. The authors of [67] state that the case $n=2m$ reduces to [67, Theorem 4.4]. In turn, the proof of [67, Theorem 4.4] relies heavily on [67, Theorem 3.5]. From Remark 5.8, we know that the proof of [67, Theorem 3.5] is incorrect.
The following example demonstrates that there exist HTM polynomials that, despite not satisfying Condition C (14), are nevertheless Hurwitz polynomials.
Example 5.10.
Let the Markov parameters, also called moments, be given by
$s_{0}=\begin{pmatrix}4&i\\ -i&4\end{pmatrix},\quad s_{1}=\begin{pmatrix}1&i\\ -i&2\end{pmatrix},\quad s_{2}=\begin{pmatrix}3&i\\ -i&4\end{pmatrix}.$
The corresponding block Hankel matrix $H_{1,1}$ and $H_{2,0}$ are positive definite matrices. According to parts (b) and (f) of Theorem 3.2, the matrix
${\bf f}_{3}(z)=$ $(-1)(Q_{2,1}^{*}(-\bar{z}^{2})+z\,P_{2,1}^{*}(-\bar{z}^{2}))$
$=$ $Iz^{3}+A_{1}z^{2}+A_{2}z+A_{3}$
is an HTM polynomial, where
$A_{1}=\left(\begin{array}[]{cc}4&i\\ -i&4\\ \end{array}\right),A_{2}=\left(\begin{array}[]{cc}5&-2i\\ 2i&3\\ \end{array}\right),A_{3}=\left(\begin{array}[]{cc}17&-6i\\ 4i&8\\ \end{array}\right).$ (140)
By using (140), we have
$A_{1}-A_{1}^{*}=0,$
$A_{3}^{*}A_{2}-A_{2}^{*}A_{3}=0,$
$A_{3}-A_{3}^{*}=\begin{pmatrix}0&-2i\\ -2i&0\\ \end{pmatrix},$ (141)
$A_{1}^{*}A_{2}-A_{2}^{*}A_{1}=\begin{pmatrix}0&-2i\\ 2i&0\\ \end{pmatrix}.$ (142)
From (141) and (142), we see that Condition C (14) is not satisfied. On the other hand, the determinant of ${\bf f}_{3}$ is given by
$\det{\bf f}_{3}=z^{6}+8z^{5}+23z^{4}+61z^{3}+121z^{2}+71z+112.$
The approximate value of its roots are
$z_{1}=$ $-4.903,z_{2}=-2.961,z_{3}=-0.055-1.117i,$
$z_{4}=$ $-0.055+1.117i,z_{5}=-0.012-2.482i,z_{6}=-0.012+2.482i.$
Consequently, ${\bf f}_{3}$ is a Hurwitz matrix polynomial.
As indicated in [12], not every Hurwitz matrix polynomial is an HTM polynomial. In the following subsection, we propose a procedure to complete a given matrix polynomial into an HTM polynomial.
5.1 Completing a matrix polynomial
In [12], the polynomial
$P_{3}(z)=I_{2}z^{3}+A_{1}z^{2}+A_{2}z+A_{3}$ (143)
with
$A_{1}=\begin{pmatrix}\frac{2187}{109}&\frac{1206}{109}\\ 0&9\end{pmatrix},\quad A_{2}=\begin{pmatrix}\frac{10662}{109}&\frac{8700}{109}\\ 0&18\end{pmatrix},\quad A_{3}=\begin{pmatrix}\frac{11178}{109}&\frac{10524}{109}\\ 0&6\end{pmatrix}$ (144)
was suggested as a polynomial that is a Hurwitz polynomial but not an HTM polynomial. The rational function $\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}$ does not satisfy property (72).
The following result allows determining whether a polynomial $P_{n}$ that is not an HTM polynomial is in fact a Hurwitz polynomial by “completing” $P_{n}$ with another polynomial $Q_{n-1}$.
Theorem 5.11.
Let $P_{n}(z)$ be $q\times q$ a monic polynomial of degree $n$. If there exists a $q\times q$ polynomial $Q_{n-1}(z)$ of degree $n-1$ such that
${\bf f}_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ (145)
is an HTM polynomial, then the polynomial $P_{n}(z)$ is a Hurwitz polynomial.
Proof.
From Theorem 3.2 f) and (75), there are polynomials $P_{1,n}$ y $Q_{1,n}$ such that $P_{n}(z)=(-1)^{n}P_{1,n}^{*}(-\bar{z})$ $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$ and $P_{1,n}(z)$ is an orthogonal polynomial on $[0,+\infty)$ with respect to a certain positive measure $\sigma$ corresponding to the Stieltjes transform $-\frac{Q_{1,n}^{*}(\bar{z})}{P_{1,n}^{*}(\bar{z})}$. By [25, Theorem 1], the roots of the polynomial $\det P_{1,n}(z)$ lie in $[0,+\infty)$. Consequently, the roots of polynomial $\det P_{n}(z)=\det((-1)^{n}P_{1,n}^{*}(-\bar{z}))$ lie in $(-\infty,0]$. ∎
Remark 5.12.
Let ${\bf f}_{n}$ be a monic matrix polynomial and ${\bf f}_{n}^{\intercal}$ be its transpose. Since $\det{\bf f}_{n}=\det{\bf f}_{n}^{\intercal}$, Theorem 5.11 is valid for $P_{n}^{\intercal}$ and $Q_{n-1}^{\intercal}$.
For finding polynomial $Q_{n-1}$ as in Theorem 5.11, one can use the following algorithm.
Algorithm for finding $Q_{n-1}$ from $P_{n}$
Step 1. Let $P_{1,n}(z)=(-1)^{n}P_{n}^{*}(-\bar{z})$ (resp. $P_{1,n}(z)=(-1)^{n}P_{n}^{\intercal^{*}}(-\bar{z})$). By using (44), compute $Q_{1,n}(z)$ in terms of the coefficients of $P_{n}(z)$ (resp. $P_{n}^{\intercal}(z)$) and the moments or Markov parameters $s_{j}$, associated with some positive measure $\sigma$ on $[0,+\infty)$.
Step 2. Verify that $\frac{P_{1,n}^{*}(\bar{z})}{Q_{1,n}^{*}(\bar{z})}$ satisfies Equality (72). If Equality (72) is not satisfied, we cannot apply this algorithm.
Step 3. By using the rational matrix function $\frac{P_{1,n}^{*}(\bar{z})}{Q_{1,n}^{*}(\bar{z})}$, calculate the remaining Markov parameters from the set $(s_{j})_{j=0}^{2n-1}$.
Step 4. If sequence $(s_{j})_{j=0}^{2n-1}$ is a positive definite Stieltjes sequence, meaning that the corresponding Hankel matrices $H_{1,n-1}$ and $H_{2,n-1}$ are positive definite matrices, then $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$.
Example 5.13.
We apply the completion procedure to the matrix polynomial (143). For $P_{3}^{\intercal}(z)$, we use the above algorithm together with Remark 5.12. With moments $s_{0}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix}$, $s_{1}=\begin{pmatrix}4b&-b\\ -b&b\end{pmatrix}$ and $s_{2}=\begin{pmatrix}16b&-2b\\ -2b&2b\end{pmatrix}$, we calculate $Q_{1,3}(z)$; therefore, we have
$Q_{1,3}(z)=s_{0}z^{2}+z(s_{1}-A_{1}^{\intercal^{*}}s_{0})+s_{2}-A_{1}^{\intercal^{*}}s_{1}+A_{2}^{\intercal^{*}}s_{0}.$ (146)
From the Laurent expansion of $-\frac{P_{1,3}^{*}(\bar{z})}{Q_{1,3}^{*}(\bar{z})}$, we calculate the moments
$s_{3}=\begin{pmatrix}96b&-6b\\ -6b&6b\end{pmatrix},s_{4}=\begin{pmatrix}768b&-24b\\ -24b&24b\end{pmatrix},s_{5}=\begin{pmatrix}7680b&-120b\\ -120b&120b\end{pmatrix}.$
The block Hankel matrices $H_{1,2}$ and $H_{2,2}$ are positive definite matrices for $b>0$. From (75), polynomial $Q_{2}(z)$ constructed from $P_{3}$ has the form $Q_{2}(z)=Q_{1,3}^{*}(-\bar{z})$. Thus, the resulting HTM polynomial ${\bf f}_{6}$ can be represented as in (145):
${\bf f}_{6}(z)=z^{6}I_{2}+B_{1}z^{5}+A_{1}z^{4}+B_{2}z^{4}+A_{2}z^{2}+B_{3}z+A_{3}.$ (147)
The coefficients $A_{1}$, $A_{2}$ and $A_{3}$ are given by (144), and the coefficients $B_{1}$, $B_{2}$ and $B_{3}$ are
$B_{1}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix},B_{2}=\begin{pmatrix}\frac{2732b}{109}&-8b\\ -8b&8b\\ \end{pmatrix},B_{3}=\begin{pmatrix}\frac{6826b}{109}&-11b\\ -11b&11b\end{pmatrix}.$ (148)
By Theorem 5.3, matrix polynomial (147) is a Hurwitz matrix polynomial for $b>0$. Finally, by Theorem 5.11, polynomial $P_{3}$ as in (143) is a Hurwitz matrix polynomial.
In the scalar version, one can construct a sequence of scalar Hurwitz polynomials $(f_{j})_{j=1}^{\infty}$ using the sequence of scalar orthogonal polynomials on $[0,+\infty)$ $(p_{1,j})_{j=1}^{\infty}$ and $(p_{2,j})_{j=1}^{\infty}$ as in Definition 2.8; see [14].
In the following lemma, we propose constructing a family of scalar Hurwitz polynomials involving the polynomials $p_{1,j}$, and $p_{2,j}$ and their derivatives.
Lemma 5.14.
For $q=1$, let $(p_{1,m}(z))$ (resp. $(p_{2,m}(z))$) be as in Definition 2.8, and let $\widetilde{p}_{1,m-1}(z):=\frac{d}{dz}p_{1,m}(z)$ (resp. $\widetilde{p}_{2,m-1}(z):=\frac{d}{dz}p_{2,m}(z)$). Thus, for $m\geq 2$, the polynomials
$f_{2m}(z)=(-1)^{m}(p_{1,m}(-z^{2})-z\widetilde{p}_{1,m-1}(-z^{2}))$ (149)
and
$f_{2m+1}(z)=(-1)^{m}(\widetilde{p}_{2,m}(-z^{2})+zp_{2,m}(-z^{2})).$ (150)
are scalar Hurwitz polynomials.
Proof.
The proof of (149) and (150) readily follows from the interlacing properties of the roots of the orthogonal polynomials $p_{k,m}(z)$ and $\widetilde{p}_{k,m-1}(z)$ for $k=1,2$, and from the fact that $a_{0}a_{1}>0$, where $a_{0}$ and $a_{1}$ are the leading coefficients of the polynomial $f_{n}(z)=a_{0}z^{n}+a_{1}z^{n-1}+\cdots+a_{n}$; see the Hermite–Biehler theorem [5], [39], [58, page 46]. ∎
Remark 5.15.
From polynomial (77), we observe that the coefficients of an HTM polynomial may be complex matrices, whereas in the scalar case the necessary condition for Hurwitzness requires all coefficients of a Hurwitz polynomial to be positive. On the other hand, from (147) and (77), we see that the determinants of the coefficients of these polynomials are positive. Based on this observation, we propose the conjecture below.
Conjecture. The determinants of the coefficients $A_{j}$ of an HTM polynomial ${\bf f}_{n}$ are positive.
Remark 5.16.
Although all known HTM polynomials appear to be Hurwitz, it remains an open question whether the Hurwitz property holds for HTM polynomials that do not satisfy Condition C (14).