### v3

# On the Hurwitz Stability of Hurwitz-Type Matrix Polynomials

Abstract
Every matrix polynomial ${\bf f}_{n}$ can be written in the form
${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2}).$
The matrix polynomial ${\bf f}_{2m}$ is said to be of Hurwitz type if the expression ${\bf g}_{2m}(z){\bf h}_{2m}^{-1}(z)$ admits a representation as a finite continued fraction with positive definite matrix coefficients. Similarly, the odd-degree matrix polynomial ${\bf f}_{2m+1}$ is of Hurwitz type if $\frac{1}{z}{\bf h}_{2m+1}(z){\bf g}_{2m+1}^{-1}(z)$ has the same property.
We derive an explicit form of the Bezoutian associated with Hurwitz-type matrix polynomials. Using this explicit form, we provide an explicit proof that Hurwitz-type matrix polynomials are Hurwitz matrix polynomials. In [52], the Hurwitzness of Hurwitz-type matrix polynomials was also studied.
Finally, we propose an extension of the class of Hurwitz-type matrix polynomials by adding to a non–Hurwitz-type matrix polynomial another matrix polynomial so that the resulting matrix polynomial is of Hurwitz type.
AMS Subject Classification (2020): 34D20 33C45 47A56 15A24 30E05
Keywords: Hurwitz matrix polynomial, orthogonal matrix polynomial, Stieltjes transform, Bezoutian, Hurwitz-type matrix polynomial
1 Introduction
A scalar polynomial with all its roots in the left half of the complex plane is called a Hurwitz polynomial [50, 35, 48, 16, 28, 34, 38, 7, 36, 37]. The Hurwitz stability of such polynomials is also referred to as Hurwitzness—a term we adopt in this work. Let $n$ and $q$ be natural numbers. In the present work, we consider the properties and Hurwitzness of a subset of Lambda matrices [45, page 95], [29, page 228], [17], defined by
${\bf f}_{n}(z):=A_{0}z^{n}+A_{1}z^{n-1}+\ldots+A_{n},$ (1)
where $A_{j}$ is a complex $q\times q$ matrix and $z$ is a complex variable. As in [41], we refer to a matrix of the form (1) as a matrix polynomial, which is called a Hurwitz matrix polynomial if $\det{\bf f}_{n}$ is a Hurwitz polynomial.
The matrix polynomial (1) has degree $n$ if $A_{0}$ is nonzero. Throughout this paper, we assume that $\det A_{0}\neq 0$. Clearly, (1) can be represented as
${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2}),$ (2)
where
${\bf h}_{n}(z)$ $:=\left\{\begin{array}[]{ll}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{array}\right$. (5)
and
${\bf g}_{n}(z)$ $:=\left\{\begin{array}[]{ll}A_{1}z^{m-1}+A_{3}z^{m-2}+\ldots+A_{2m-1},&n=2m,\\ A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m+1.\end{array}\right$. (8)
We will refer to (5) as the ${\bf h}_{n}$ part and to (8) as the ${\bf g}_{n}$ part of the matrix polynomial ${\bf f}_{n}$.
Let us introduce the following notation. The symbol ${\mathbb{C}}^{q\times q}$ represents the set of all complex $q\times q$ matrices. For the null matrix that belongs to ${\mathbb{C}}^{p\times q}$, we will write $0_{p\times q}$. We denote by $0_{q}$ and $I_{q}$ the null and identity matrices in ${\mathbb{C}}^{q\times q}$, respectively. When the sizes of the null and identity matrices are clear from the context, we will omit their indices.
In the following definition, we recall the concept of a Hurwitz-type matrix (HTM) polynomial [5]. In [5], HTM polynomials were referred to as matrix Hurwitz-type polynomials. For $A,\,B\in{\mathbb{C}}^{q\times q}$ with $B$ invertible, set $\frac{A}{B}:=AB^{-1}$.
Definition 1.1.
The $q\times q$ matrix polynomial ${\bf f}_{n}$ in (1) is called a Hurwitz-type matrix polynomial if there exist two sequences of positive definite matrices, $({\bf c}_{k})_{k=0}^{m-1}$ and $({\bf d}_{k})_{k=0}^{m-1}$ (resp. $({\bf c}_{k})_{k=0}^{m-1}$ and $({\bf d}_{k})_{k=0}^{m}$) such that, for $n=2m$,
$\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}=$ $\cfrac{I_{q}}{z{\bf c}_{0}+\cfrac{I_{q}}{{\bf d}_{0}+\cfrac{I_{q}}{\ddots+{\bf d}_{m-2}+\cfrac{I_{q}}{+z{\bf c}_{m-1}+{\bf d}_{m-1}^{-1}}}}},$ (9)
for all $z\in{\mathbb{C}}$ with $\det{\bf h}_{n}(z)\neq 0$ and, for $n=2m+1$,
$\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}=$ $\cfrac{I_{q}}{z{\bf c}_{0}+\cfrac{I_{q}}{{\bf d}_{0}+\cfrac{I_{q}}{\ddots\,+z{\bf c}_{m-1}+\cfrac{I_{q}}{{\bf d}_{m-1}+z^{-1}{\bf c}_{m}^{-1}}}}}$ (10)
for all $z\in{\mathbb{C}}\setminus\{0\}$ with $\det{\bf g}_{n}(z)\neq 0$.
The definition of HTM polynomials in the form of a finite continued fraction is inspired by [30, Chapter XV, Theorem 15].
Main Results. Below are the main results of the present work:
  • a)
    We compute explicitly the Bezoutian of the HTM polynomial (1); see Theorem 3.4 and Corollary 3.5. From this, an explicit proof of the Hurwitz property of HTM polynomials readily follows; see Theorem 4.3.
  • b)
    For a given matrix polynomial $P_{n}$ of degree $n$ that is not a Hurwitz-type matrix (HTM) polynomial, we construct a polynomial ${\bf f}_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ of degree $2n$, with a suitably chosen $Q_{n-1}$, such that ${\bf f}_{2n}$ becomes an HTM polynomial. Then, by applying the result from [22], we conclude that $P_{n}$ itself is an HTM polynomial; see Theorem 5.1.
The Hurwitz property of matrix polynomials via Bezoutians was studied in [31, 43, 44]. Note that in [52] the Hurwitz property of HTM polynomials was examined, but the proof of [52, Theorem 3.5], which serves as the basis for the analysis of the Hurwitz property of HTM polynomials in [52], is not fully detailed. More precisely, the proof of [52, Theorem 3.5] leaves certain steps implicit; see Remark 6.2. In particular, the argument is developed only for the case $n=2m$, whereas the case $n=2m+1$ is referred to as “similar,” without explicit justification.
Additionally, in contrast to [52], we use representations of the matrix polynomial ${\bf f}_{n}$ in terms of orthogonal polynomials on $0,\infty)$ of the first kind and polynomials of the second kind, as in (2) and (24), for example, to compute explicitly the Bezoutian of the quadruple $({\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x))$.
The asterisk denotes the Hermitian adjoint of a matrix, while the overbar represents complex conjugation. In the scalar case, the use of the ${\bf h}_{n}$ part and ${\bf g}_{n}$ part of the polynomial ${\bf f}_{n}$ for studying the corresponding Bezoutian was employed by Krein and Naimark in [40] and by Krein and Akhiezer in [1].
The importance of HTM polynomials as a subset of Hurwitz matrix polynomials is that they help study the asymptotic stability of the differential equation
$I_{q}y^{(n)}(t)+A_{1}y^{(n-1)}(t)+\ldots+A_{n-1}y^{\prime}(t)+A_{m}y(t)=0,\ \ t\geq 0$
with $y(t)\in{\mathbb{R}}^{n}$; $I_{q},A_{1},\ldots,A_{n}\in{\mathbb{C}}^{q\times q}$; see [46, 51, 49, 32] and [39].
Our motivation is to generalize the results regarding robust stability [6] and finite-time stabilization with bounded controls [12, 13, 8, 10] for the matrix case within the framework of HTM polynomials.
Our work is organized as follows. In Section 2, we review previous results regarding HTM polynomials and their associated Markov parameters. Moreover, we mention Stieltjes orthogonal matrix polynomials, orthogonal matrix polynomials, interrelations between HTM polynomials and Markov parameters, the Bezoutian, and the inertia of matrix polynomials.
In Section 3, we explicitly compute the Bezoutian of HTM polynomials. In Section 4, we study the Hurwitzness of HTM polynomials. In Section 5, we present a result and an algorithm for constructing an HTM polynomial from a given matrix polynomial that is not an HTM polynomial. Additionally, we indicate why the original matrix polynomial is a Hurwitz matrix polynomial. In Section 6, we explain in more detail certain steps left implicit in Theorem 3.5 of [52] on a “matrix analogue of the stability criterion via Markov parameters,” as stated on page 10 of [52].
This work is a direct continuation of the author’s research in [4] and [5], which focus on the nondegenerate truncated matricial Stieltjes problem and the HTM polynomials, respectively.
2 Preliminaries
In this section, we present the notions and results that will be used throughout this work.
2.1 HTM polynomials and associated Markov parameters
In the following remark, we reproduce the Laurent expansion of $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$, which allows us to relate the matrix polynomial (1) with the so-called Markov parameters $s_{j}$ of the polynomial ${\bf f}_{n}$.
Remark 2.1.
Let ${\bf h}_{n}$ and ${\bf g}_{n}$ be defined as in (5) and (8), and let ${\bf f}_{n}$ be as in (2). Let $|z_{0}|\geq\max\{|\lambda|:\det{\bf h}_{n}(\lambda)=0\}$ (resp. $|z_{0}|\geq\max\{|\lambda|:\det{\bf g}_{n}(\lambda)=0\}$). For $|z|>|z_{0}|$, the power series expansion of ${\bf g}_{n}/{\bf h}_{n}$ (resp. ${\bf h}_{n}/(z{\bf g}_{n}$)) in negative powers of $z$ yields
$\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ $=\frac{s_{0}}{z}-\frac{s_{1}}{z^{2}}+\ldots+(-1)^{n}\frac{s_{n}}{z^{n+1}}+\ldots\,,\ \ n=2m,$ (11)
$\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ $=\frac{s_{0}}{z}-\frac{s_{1}}{z^{2}}+\ldots+(-1)^{n}\frac{s_{n}}{z^{n+1}}+\ldots\,,\ \ n=2m+1.$ (12)
For $j\geq 0$ and by using the Markov parameters, we construct block Hankel matrices:
$H_{1,j}:=\left(\begin{array}[]{cccc}s_{0}&s_{1}&\ldots&s_{j}\\ s_{1}&s_{2}&\ldots&s_{j+1}\\ \vdots&\vdots&\vdots&\vdots\\ s_{j}&s_{j+1}&\ldots&s_{2j}\end{array}\right),H_{2,j}:=\left(\begin{array}[]{cccc}s_{1}&s_{2}&\ldots&s_{j+1}\\ s_{2}&s_{3}&\ldots&s_{j+2}\\ \vdots&\vdots&\vdots&\vdots\\ s_{j+1}&s_{j+2}&\ldots&s_{2j+1}\end{array}\right).$ (13)
The rational matrix functions in (9) and (10) are special solutions to the truncated Stieltjes matrix moment (TSMM) problem; see proof of [5, Theorem 7.9]. This implies that $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ are Stieltjes transforms of a certain positive measure $\sigma$ on $0,+\infty)$.
The TSMM problem is defined as follows: Let $(s_{j})_{j=0}^{n}$ be a finite sequence of $q\times q$ complex Hermitian matrices. The TSMM problem consists of finding the set ${\mathscr{M}}_{n}$ of all nonnegative Hermitian $q\times q$ measures $\sigma$ defined on the Borel $\sigma$–algebra ${\mathfrak{B}}\cap[0,+\infty)$ such that
$s_{j}=\int\limits_{[0,+\infty)}t^{j}\sigma(dt),\quad 0\leq j\leq n-1,\quad s_{n}=\int\limits_{[0,+\infty)}t^{n}\sigma(dt)+M,$
where $M$ is a $q\times q$ complex valued positive semi-definite matrix.
For each $\sigma\in{\mathscr{M}}_{n}$, one typically associates it with a matrix function
$s(z)=\int_{[0,+\infty)}\frac{\sigma(dt)}{t-z},$
which is defined and holomorphic in ${\mathbb{C}}\setminus[0,+\infty)$. The matrix function $s$ is said to be associated with the TSMM problem. See [18, 19, 20, 21, 23, 24, 14, 25], [26, 27].
In [18], it is proven that the TSMM problem in the case of an odd (resp. even) number of moments has a solution if and only if $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are both positive semi-definite matrices.
Definition 2.2.
The sequence $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is called a Stieltjes positive sequence if both corresponding matrices $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are positive definite. In the case where $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is a positive Stieltjes sequence, the corresponding TSMM problem is called a TSMM problem in the nondegenerate case.
In [18] and [19], Yu. Dyukarev found the complete set of solutions to the TSMM problem in the case that $(s_{j})_{j=0}^{2m}$ and $(s_{j})_{j=0}^{2m+1}$ are Stieltjes positive sequences, respectively.
2.2 Stieltjes orthogonal matrix polynomials
We now reproduce from [4] a number of auxiliary matrices, as well as two families of orthogonal matrix polynomials and their polynomials of the second kind. Let $R_{j}:\mathbb{C}\to\mathbb{C}^{(j+1)q\times(j+1)q}$ be given by
$R_{j}(z):=(I_{(j+1)q}-zT_{j})^{-1},\qquad j\geq 0,$
with
$T_{0}:=0_{q},\ \ T_{j}:=\left(\begin{array}[]{cc}0_{q\times jq}&0_{q}\\ I_{jq}&0_{jq\times q}\\ \end{array}\right),\qquad j\geq 1.$
Observe that for each $j\in{\mathbb{N}}_{0}$, the matrix–valued function $R_{j}$ can be represented via
$R_{j}(z)=\left(\begin{array}[]{cccccc}I_{q}&0_{q}&0_{q}&\ldots&0_{q}&0_{q}\\ zI_{q}&I_{q}&0_{q}&\ldots&0_{q}&0_{q}\\ z^{2}I_{q}&zI_{q}&I_{q}&\ldots&0_{q}&0_{q}\\ \vdots&\vdots&\vdots&\iddots&\vdots&\vdots\\ z^{j}I_{q}&z^{j-1}I_{q}&z^{j-2}I_{q}&\ldots&zI_{q}&I_{q}\end{array}\right).$ (14)
Let
$v_{0}:=$ $I_{q},\quad v_{j}:=\left(\begin{array}[]{c}I_{q}\\ 0_{jq\times q}\\ \end{array}\right)\ \,\mbox{for all}\ \,j\in\mathbb{N}.$ (17)
Furthermore, let
$Y_{1,j}:=y_{[j,2j-1]},\,\ 1\leq j\leq n,\,\quad\mbox{and}\quad Y_{2,j}:=y_{[j+1,2j]},\,\ 2\leq j\leq n,$
with
$y_{[j,k]}:=\left(s_{j},s_{j+1},\ldots,s_{k}\right)^{*},\quad 0\leq j\leq k$
and $y_{[j,k]}=0_{q}$ if $j>k$. Define
$u_{1,0}:=0_{q},\quad u_{1,j}:=\left(\begin{array}[]{c}0_{q}\\ -y_{[0,j-1]}\\ \end{array}\right),\quad u_{2,j}:=\begin{array}[]{c}-y_{[0,j]}\\ \end{array},\quad 1\leq j\leq n.$ (21)
Let $\widehat{H}_{1,j}$ (resp. $\widehat{H}_{2,j}$) denote the Schur complement of the block Hankel matrices $H_{1,j-1}$ in $H_{1,j}$ (resp.  $H_{2,j-1}$ in $H_{2,j}$) as seen in [47], [4, Equalities (2.5) and (2.6)]:
$\widehat{H}_{1,0}:=s_{0},\quad\widehat{H}_{1,j}:=$ $s_{2j}-Y_{1,j}^{*}H_{1,j-1}^{-1}Y_{1,j},\ \ j\geq 1,$
$\widehat{H}_{2,0}:=s_{1},\quad\widehat{H}_{2,j}:=$ $s_{2j+1}-Y_{2,j}^{*}H_{2,j-1}^{-1}Y_{2,j},\ \ j\geq 1.$
These matrices are Hermitian positive definite matrices, and so are $H_{1,j}$ and $H_{2,j}$.
Now we partially reproduce left orthogonal matrix polynomials (OMP) on $0,+\infty) [$22].
Definition 2.3.
Let $\sigma$ be a $q\times q$ nonnegenative measure on $0,+\infty)$. A sequence $(P_{j})_{j=0}^{\infty}$ of $q\times q$ matrix-valued polynomials is called orthogonal with respect to $\sigma$ if
each matrix-valued polynomial $P_{j}$ is of degree $j$;
the matrix-valued polynomials $P_{j}$ satisfy the orthogonality relations:
$\int\limits_{[0,+\infty)}P_{j}(t)\,\sigma(dt)\,P_{k}^{*}(t)=\delta_{jk}C_{qj},$
for all $j,k\in\mathbb{N}$. Here, $\delta_{jk}$ is the Kronecker delta, and and $C_{qj}$ is a constant nonnegenative $q\times q$ matrix.
The following matrix polynomials seem to have been first introduced in [21].
Definition 2.4.
Let $(s_{j})_{j=0}^{2n}$ (resp. $(s_{j})_{j=0}^{2n+1}$) be a Stieltjes positive sequence. Furthermore, let $v_{j}$, $u_{1,j}$, $u_{2,j}$ and $R_{j}$ be as in (17), (21) and (14), respectively. Let
$P_{1,0}(z):=I_{q},\ \ Q_{1,0}(z):=0_{q},\ \ P_{2,0}(z):=I_{q},\ \ Q_{2,0}(z):=s_{0}.$
For $j\geq 1$, set
$P_{1,j}(z):=$ $(-Y_{1,j}^{*}H_{1,j-1}^{-1},\,I_{q})R_{j}(z)v_{j},$
$P_{2,j}(z):=$ $(-Y_{2,j}^{*}H_{2,j-1}^{-1},\,I_{q})R_{j}(z)v_{j},$
$Q_{1,j}(z):=$ $-(-Y_{1,j}^{*}H_{1,j-1}^{-1},\,I_{q})R_{j}(z)u_{1,j},$
$Q_{2,j}(z):=$ $-(-Y_{2,j}^{*}H_{2,j-1}^{-1},\,I_{q})R_{j}(z)u_{2,j}.$
The matrix polynomials $Q_{1,j}$ and $Q_{2,j}$ are called second-kind polynomials with respect to $P_{1,j}$ and $P_{2,j}$, respectively.
In the next proposition, we recall that the matrix polynomials $P_{1,j}$ and $P_{2,j}$ are orthogonal with respect to a positive measure supported on $0,+\infty)$, and that the associated second-kind polynomials $Q_{1,j}$ and $Q_{2,j}$ can be obtained from $P_{1,j}$ and $P_{2,j}$, respectively. We also record an identity involving $P_{k,j}$ and $Q_{k,j}$ whose value is the zero matrix.
Proposition 2.5.
a) The polynomials $P_{1,j}$ and $P_{2,j}$ are OMP with respect to $\sigma$ and $t\sigma$, respectively. More precisely (as seen in [14, Remark 6.3]),
$\int\limits_{[0,+\infty)}P_{k,j}(t)t^{k-1}\sigma(dt)P_{k,l}^{*}(t)=\left\{\begin{array}[]{ccc}0_{q}&:&j\neq l\\ \widehat{H}_{k,j}&:&j=l\end{array}\right.,\ \ k=1,2.$
b) The following identities hold (as seen in [4, Remark D.6] and [4, Remark E.4]):
$Q_{1,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{P_{1,j}(x)-P_{1,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n,$ (22)
$Q_{2,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{xP_{2,j}(x)-tP_{2,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n-1.$
c) For $k=1,2$, the following identity holds (as seen in [4, Proposition 5.1 (a)])
$P_{k,j}(z)Q_{k,j}^{*}(\bar{z})-Q_{k,j}(z)P_{k,j}^{*}(\bar{z})=$ $0_{q}.$ (23)
For $n\geq 1$, it follows from [5, Theorem 6.1] that every HTM polynomial admits the representation
${\bf f}_{n}(z)=\left\{\begin{array}[]{cl}(-1)^{m}(P_{1,m}^{*}(-\bar{z}^{2})-z\,Q_{1,m}^{*}(-\bar{z}^{2})),&n=2m,\\ (-1)^{m}(Q_{2,m}^{*}(-\bar{z}^{2})+z\,P_{2,m}^{*}(-\bar{z}^{2})),&n=2m+1,\end{array}\right$. (24)
Identity (24) is obtained by using the relations
${\bf h}_{n}(z)=$ $(-1)^{m}P_{1,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m+1}Q_{1,m}^{*}(-\bar{z}),\qquad n=2m$ (25)
and
${\bf h}_{n}(z)=$ $(-1)^{m}Q_{2,m}^{*}(-\bar{z})\quad\mbox{and}\quad{\bf g}_{n}(z)=(-1)^{m}P_{2,m}^{*}(-\bar{z}),\qquad n=2m+1$ (26)
which are established in the proof of [5, Proposition 7.8].
The notation below was introduced in [5, Lemma 2.6]. If $n=2m$, we set ${\mathscr{N}}_{{\bf h}_{n}}:=\{z\in{\mathbb{C}}:\,\det{\bf h}_{n}(z)=0\}$ , whereas for $n=2m+1$ we set ${\mathscr{N}}_{{\bf g}_{n}}:=\{z\in{\mathbb{C}}:\,\det{\bf g}_{n}(z)=0\}$. Furthermore, we define $\eta_{n}:=\max\{|z|:\,z\in{\mathscr{N}}_{{\bf h}_{n}}\}$, respectively, $\rho_{n}:=\max\{|z|:\,z\in{\mathscr{N}}_{{\bf g}_{n}}\}$. In the proof of [5, Theorem 6.1], the so–called extremal solutions to the TSMM problem are expressed as
$G_{\sigma_{2m,\max}}(z)=-\frac{Q_{1,j}^{*}(\bar{z})}{P_{1,j}^{*}(\bar{z})}=\frac{{\bf g}_{2m}(-z)}{{\bf h}_{2m}(-z)}$ (27)
and
$G_{\sigma_{2m,\min}}(z)=-\frac{Q_{2,j}^{*}(\bar{z})}{zP_{2,j}^{*}(\bar{z})}=\frac{{\bf h}_{2m+1}(-z)}{(-z){\bf g}_{2m+1}(-z)}.$ (28)
Note that the matrix function $G_{\sigma_{2m,\max}}$ (resp. $G_{\sigma_{2m,\min}}$) is defined for all $z\in{\mathbb{C}}\setminus[0,\infty)$ with $|z|>\eta_{2m}$ (resp. for all $z\in{\mathbb{C}}\setminus[0,\infty)$ with $|z|>\rho_{2m+1}$).
2.3 The Bezoutian and inertia for matrix polynomials
In this subsection, we reproduce some notions related to the inertia for a matrix polynomial [43, 44]. The inertia for scalar polynomials was considered in [42, page 462].
Let $L(\lambda)=L_{0}\lambda^{\ell}+L_{1}\lambda^{\ell-1}+\ldots+\lambda L_{\ell-1}+L_{\ell}$ be a $q\times q$ matrix polynomial with coefficients $L_{k}\in{\mathbb{C}}^{q\times q}$. If $L_{0}\neq 0_{q}$, then ${\rm deg}(L)=\ell$ is the degree of polynomial $L$. Furthermore, if $L_{0}=I_{q}$, then $L$ is called monic. Let $M_{1},L_{1},M$ and $L$ be polynomial matrices such that $M_{1}(\lambda)L_{1}(\lambda)-M(\lambda)L(\lambda)=0$. The Bezoutian associated with the quadruple $(M_{1},L_{1},M,L)$ is the block matrix
$B_{M_{1},M}(L,L_{1})=\left(\Gamma_{j,k}\right)_{j,k=1}^{m-1,\ell-1}.$ (29)
The block entries $\Gamma_{j,k}$ of the quadruple $(M_{1},L_{1},M,L)$ are determined by following equality:
$\sum_{j,k=0}^{m=1,\ell-1}\Gamma_{jk}\lambda^{j}\mu^{k}=\frac{1}{\lambda-\mu}\left(M_{1}(\lambda)L_{1}(\mu)-M(\lambda)L(\mu)\right).$ (30)
The quantity $\ell$ (resp. $m$) denotes the maximal degree of $L(\lambda)$, $L_{1}(\lambda)$ (resp. $M(\lambda)$, $M_{1}(\lambda)$). Assume that there exists $\lambda_{0}\in{\mathbb{C}}$ such that $\det L(\lambda_{0})\neq 0$: on this case $L$ is called regular.
Let $L$ be a regular polynomial of degree $\ell$. The triple
$\gamma(L)=\left(\gamma_{+}(L),\gamma_{-}(L),\gamma_{0}(L)\right)$ (31)
is called inertia with respect to the real axis, where $\gamma_{+}(L)$ (resp. $\gamma_{-}(L)$) is the number of zeros of $\det(L(\lambda))$ in the open upper (resp. open lower) half-plane, and $\gamma_{0}(L)$ is the number of zeros of $\det(L(\lambda))$ in the real axis. Now we reproduce [43, Corollary 2.2].
Corollary 2.6.
Let $L(\lambda)$ be a matrix polynomial with an invertible leading coefficient. The spectrum of $L(\lambda)$ lies in the upper half-plane if and only if there exists a regular matrix polynomial $L_{1}(\lambda)$ that satisfies
$L_{1}^{*}(\bar{\lambda})L_{1}(\lambda)=L^{*}(\bar{\lambda})L(\lambda)$ (32)
such that the matrix
$\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})$} (33)
is positive definite.
To compute the matrix (33), we use the generalized Bezoutian matrix and form introduced by Anderson and Jury in [2], which is a $q\times q$ rational function $W(z)$. Suppose that $W(z)$ admits two matrix factorizations:
$W(z)=A^{-1}(z)B(z)=D(z)C^{-1}(z).$ (34)
The generalized Bezoutian form associated with the quadruple A,B,C,D{A,B,C,D} is defined as
$\Gamma(x,y)=\frac{1}{x-y}\left(A(x)D(y)-B(x)G(y)\right).$
Note that in this work the relation (32) and the matrix (33) are associated with the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$), while the decomposition (34) is connected to the identity
$s_{{\bf f}_{n}}(z)=s_{{\bf f}_{n}}^{*}(\bar{z}),$ (35)
where $s_{{\bf f}_{n}}(z)$ denotes the left-hand side of (9), (10).
In the following remark, we reproduce an assertion from [2, Remark 2.2] concerning the integrability of the rational function $\Gamma$ in (34): the conditions under which $\Gamma$ is a polynomial in both $x$ and $y$.
Remark 2.7.
For $\Gamma(x,y)$ to be a polynomial in both $x$ and $y$, it is necessary and sufficient that
$A(x)D(x)=B(x)C(x).$
Our strategy for analyzing the Hurwitzness of the HTM polynomials is as follows: we explicitly compute the matrix (33) by using the representations given in (80) and (81) for the quadruple $({\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x))$. Rather than applying formula (36) directly, we employ Bezoutian forms associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$, corresponding to the cases $n=2m+1$ and $n=2m$, respectively. See Equalities (37) and (38).
3 Bezoutian and Hurwitz-type matrix polynomials
In this section, using the so-called matrix symmetrizers (55) and block Hanckel matrices $H_{1,j}$ and $H_{2,j}$, we derive an explicit form of the Bezoutian for the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$).
Instead of working directly with the Bezoutian form
${\mathscr{F}}_{n}(x,y):=$ $\frac{{\bf f}_{n}^{*}(\bar{x}){\bf f}_{n}(-y)-{\bf f}_{n}^{*}(-\bar{x}){\bf f}_{n}(y)}{x-y},$ (36)
we use the Equality (2) and calculate the Bezoutian forms
${\mathscr{G}}_{n}^{(1)}(x,y):=$ $\frac{1}{x^{2}-y^{2}}\left(x^{2}{\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-y^{2}{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})\right)$ (37)
and
${\mathscr{G}}_{n}^{(2)}(x,y):=$ $\frac{xy}{x^{2}-y^{2}}\left({\bf g}_{n}^{*}(\bar{x}^{2}){\bf h}_{n}(y^{2})-{\bf h}_{n}^{*}(\bar{x}^{2}){\bf g}_{n}(y^{2})\right)$ (38)
associated with the quadruples $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$. Recall that a similar decomposition of the Bezoutian for scalar polynomials was previously studied; see [40, page 293] and [1, page 257]. Following [2] to compute the Bezoutian forms ${\mathscr{G}}_{n}^{(1)}(x,y)$ and ${\mathscr{G}}_{n}^{(2)}(x,y)$ we employ both representations (35) of the rational matrices (9) and (10).
It readily follows that
${\mathscr{F}}_{n}(x,y)=2\left({\mathscr{G}}_{n}^{(1)}(x,y)+{\mathscr{G}}_{n}^{(2)}(x,y)\right).$ (39)
In the next lemma, we demonstrate that the matrix functions ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are matrix polynomials in $x$ and $y$. See [2, Remark 1].
Lemma 3.1.
Let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as defined in (36), (37) and (38), respectively. The functions ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are polynomials on $x$ and $y$.
Proof.
Using (24) and (23), we have
${\bf f}_{n}(\bar{x}){\bf f}_{n}(-x)-{\bf f}_{n}(-\bar{x}){\bf f}_{n}(x)$
$=2x\left\{\begin{array}[]{lc}P_{1,m}(-x^{2})Q_{1,m}^{*}(-x^{2})-Q_{1,m}(-x^{2})P_{1,m}^{*}(-x^{2}),&n=2m\\ P_{2,m}(-x^{2})Q_{2,m}^{*}(-x^{2})-Q_{2,m}(-x^{2})P_{2,m}^{*}(-x^{2}),&n=2m+1\end{array}\right$.
$=0.$
By [15, Chapter 1], any polynomial vanishing on the diagonal $x=y$ is divisible by xyx-y. Since ${\bf f}_{n}^{*}(\bar{x}){\bf f}_{n}(-y)-{\bf f}_{n}^{*}(-\bar{x}){\bf f}_{n}(y)$ has this property, ${\mathscr{F}}_{n}(x,y)$ is a polynomial. Arguing similarly with (24), (23), and [15, Chapter 1], we show that ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ are polynomials in $x$ and $y$. ∎
For $k=0,1$, let
$H^{(k)}_{\infty}:=$ $\begin{pmatrix}s_{k}&-s_{k+1}&s_{k+2}&\ldots\\ -s_{k+1}&s_{k+2}&\iddots\\ s_{k+2}&\iddots\end{pmatrix}$ (40)
be an infinite Hankel matrix.
A similar untruncated block Hankel matrix, denoted by $\bf{H}$, appears in the proof of [2, Lemma 2.3].
Lemma 3.2.
Let ${\mathscr{F}}_{n}$, ${\mathscr{G}}_{n}^{(1)}$ and ${\mathscr{G}}_{n}^{(2)}$ be as in (36), (37) and (38), respectively. Thus, for $n=2m$, we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-2}I_{q},\bar{y}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}),$ (41)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(0)}_{\infty}\left(\bar{y}^{-1}I_{q},\bar{y}^{-3}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}).$ (42)
For $n=2m+1$, we have
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ ${\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(I_{q},x^{-2}I_{q},\ldots\right)H^{(0)}_{\infty}\left(I_{q},\bar{y}^{-2}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(y^{2}),$ (43)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-{\bf g}_{2m+1}^{*}(\bar{x}^{2})\left(x^{-1}I_{q},x^{-3}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-1}I_{q},\bar{y}^{-3}I_{q},\ldots\right)^{*}$
$\cdot{\bf g}_{2m+1}(y^{2}).$ (44)
Proof.
We prove (41). With (11), we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ $\frac{1}{x^{2}-y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{2}{\bf h}_{2m}^{-1^{*}}(\bar{x}^{2}){\bf g}_{2m}^{*}(\bar{x}^{2})-y^{2}{\bf g}_{2m}(y^{2}){\bf h}_{2m}^{-1}(y^{2})\right)$
$\cdot{\bf h}_{2m}(y^{2})$
$=$ $\frac{1}{x^{2}-y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(s_{1}(y^{-2}-x^{-2})-s_{2}(y^{-4}-x^{-4})+s_{3}(y^{-6}-x^{-6})\right$.
$\left.-\ldots\right){\bf h}_{2m}(y^{2})$
$=$ $\frac{1}{x^{2}y^{2}}{\bf h}_{2m}^{*}(\bar{x}^{2})\left(s_{1}-s_{2}(x^{-2}+y^{-2})+s_{3}(x^{-4}+x^{-2}y^{-2}+y^{-4})\right$.
$\left.-\ldots\right){\bf h}_{2m}(y^{2})$
$=$ ${\bf h}_{2m}^{*}(\bar{x}^{2})\left(x^{-2}I_{q},x^{-4}I_{q},\ldots\right)H^{(1)}_{\infty}\left(\bar{y}^{-2}I_{q},\bar{y}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(y^{2}).$
In the last equality, we substituted (40) for $k=1$. Similarly, to prove (42), we use (11) and (40) for $k=0$. Equalities (43) and (44) are proven by applying (12) and (40) for $k=0$ and $k=1$, respectively. ∎
Let
${\bf h}_{2m}^{(11)}(\xi)+{\bf h}_{2m}^{(12)}(\xi)=$ $\left(\bar{\xi}^{-2}I_{q},\bar{\xi}^{-4}I_{q},\ldots\right)^{*}{\bf h}_{2m}(\xi^{2}),$ (45)
${\bf h}_{2m}^{(21)}(\xi)+{\bf h}_{2m}^{(22)}(\xi)=$ $\left(\bar{\xi}^{-1}I_{q},\bar{\xi}^{-3}I_{q},\ldots\right)^{*}{\bf h}_{2m}(\xi^{2}),$ (46)
${\bf g}_{2m+1}^{(11)}(\xi)+{\bf g}_{2m+1}^{(12)}(\xi)=$ $\left(I_{q},\bar{\xi}^{-2}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(\xi^{2}),$ (47)
${\bf g}_{2m+1}^{(21)}(\xi)+{\bf g}_{2m+1}^{(22)}(\xi)=$ $\left(\bar{\xi}^{-1}I_{q},\bar{\xi}^{-3}I_{q},\ldots\right)^{*}{\bf g}_{2m+1}(\xi^{2}),$ (48)
where ${\bf h}_{2m}^{(k1)}(\xi)$ (resp. ${\bf g}_{2m+1}^{(k1)}(\xi)$) for $k=1,2$ are polynomials in $\xi$, and ${\bf h}_{2m}^{(k2)}(\xi)$ (resp. ${\bf g}_{2m+1}^{(k2)}(\xi)$) for $k=1,2$ are polynomials in $\xi^{-1}$.
For a natural number $m$, let
$\widetilde{J}_{m}:={\rm diag}\,\left(I_{q},(-1)I_{q},(-1)^{2}I_{q}\ldots,(-1)^{m}I_{q}\right)$ (49)
be a $(m+1)q\times(m+1)q$. The following remark readily follows from direct calculation.
Remark 3.3.
Let ${\bf h}_{2m}^{(11)}(\xi)$, ${\bf h}_{2m}^{(21)}(\xi)$, ${\bf g}_{2m+1}^{(11)}(\xi)$ and ${\bf g}_{2m+1}^{(21)}(\xi)$ be as defined in (45), (46), (47) and (48), respectively. Furthermore, let $H_{1,j}$, $H_{2,j}$ and $J_{m}^{(k)}$ be as defined in (13) and (49), respectively. Thus, the following equalities hold
${\bf h}_{2m}^{(11)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf h}_{2m}^{(11)}(y)=$ $\widehat{\bf h}_{2m}^{(11)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(11)}(y),$ (50)
${\bf h}_{2m}^{(21)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf h}_{2m}^{(21)}(y)=$ $\widehat{\bf h}_{2m}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(21)}(y),$ (51)
${\bf g}_{2m+1}^{(11)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf g}_{2m+1}^{(11)}(y)=$ $\widehat{\bf g}_{2m+1}^{(11)^{*}}(\bar{x})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}\,\widehat{\bf g}_{2m+1}^{(11)}(y),$ (52)
${\bf g}_{2m+1}^{(21)^{*}}(\bar{x})H^{(1)}_{\infty}{\bf g}_{2m+1}^{(21)}(y)=$ $\widehat{\bf g}_{2m+1}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf g}_{2m+1}^{(21)}(y),$ (53)
where
$\widehat{\bf h}_{2m}^{(11)},\,\widehat{\bf h}_{2m}^{(21)},\,\widehat{\bf g}_{2m+1}^{(11)}\quad\mbox{and}\quad\widehat{\bf g}_{2m+1}^{(11)}$
represents finite-dimensional restrictions of the matrices ${\bf h}_{2m}^{(11)}$, ${\bf h}_{2m}^{(21)}$, ${\bf g}_{2m+1}^{(11)}$ and ${\bf g}_{2m+1}^{(11)}$. In particular, the polynomial part of
$\widehat{\bf h}_{2m}^{(11)}(y)=\left(\bar{y}^{-2}{\bf h}_{2m}^{*}(\bar{y}^{2}),\bar{y}^{-4}{\bf h}_{2m}^{*}(\bar{y}^{2}),\ldots,\bar{y}^{-2m}{\bf h}_{2m}^{*}(\bar{y}^{2})\right)^{*}$ (54)
is the restriction of the polynomial part of the matrix
${\bf h}_{2m}^{(11)}(y)=\left(\bar{y}^{-2}{\bf h}_{2m}^{*}(\bar{y}^{2}),\bar{y}^{-4}{\bf h}_{2m}^{*}(\bar{y}^{2}),\ldots,\bar{y}^{-2m}{\bf h}_{2m}^{*}(\bar{y}^{2}),0_{q},0_{q},\ldots\right)^{*}.$
Equalities (50)–(53) correspond to those in the proof of [2, Lemma 2.3], where, instead of the matrices $H_{1,j}$ and $H_{2,j}$, the truncated block Hankel matrix ${\bf H}_{nm}$ is used. Moreover, in place of the matrix $H_{\infty}^{(1)}$, the infinite Hankel matrix $H$ is employed.
For a matrix polynomial $Q(x)=Q_{0}x^{\ell}+Q_{1}x^{\ell-1}+\ldots+Q_{\ell-1}x+Q_{\ell}$, the Hankel block matrix
$S(Q):=\begin{pmatrix}Q_{\ell-1}&Q_{\ell-2}&\ldots&Q_{1}&Q_{0}\\ Q_{\ell-2}&Q_{\ell-3}&\ldots&Q_{0}&0_{q}\\ \vdots&\iddots&\iddots&\iddots&\vdots\\ Q_{0}&0_{q}&\ldots&0_{q}&0_{q}\end{pmatrix}$ (55)
is called a matrix symmetrizer of $Q(x)$. In [42, page 455], the concept of symmetrizer is introduced for scalar polynomials. Additionally, let
$F_{0}^{2k}(x):=$ $\left(I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2k}I_{q}\right)^{*},$ (56)
$F_{1}^{2k-1}(x):=$ $\left(\bar{x}I_{q},\bar{x}^{3}I_{q},\ldots,\bar{x}^{2k-1}I_{q}\right)^{*}.$ (57)
In the following theorem, added by symmetrizers of corresponding matrices and the block Hankel matrices $H_{k,j}$, we obtain a factorized expression of the Bezoutian forms ${\mathscr{G}}_{2m}^{(1)}$, ${\mathscr{G}}_{2m}^{(2)}$, ${\mathscr{G}}_{2m+1}^{(1)}$ and ${\mathscr{G}}_{2m+1}^{(2)}$.
Theorem 3.4.
Let ${\mathscr{G}}_{2m}^{(k)}$, ${\mathscr{G}}_{2m+1}^{(k)}$ for $k=1,2$ be as defined in (37) and (38). Additionally, let ${\bf h}_{n}$, ${\bf g}_{n}$, $S$, $\widetilde{J}_{m}$, $H_{1,j}$, $H_{2,j}$, $F_{0}^{2k}(x)$ and $F_{1}^{2k-1}(x)$ be as in (5), (8), (55), (49), (13), (56) and (57), respectively.
Therefore, for $n=2m$, the following equality holds:
${\mathscr{G}}_{2m}^{(1)}(x,y)=$ $F_{0}^{(2m-2)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{0}^{(2m-2)}(y),$ (58)
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m})F_{1}^{(2m-1)}(y).$ (59)
For $n=2m+1$, we have
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ $F_{0}^{(2m)^{*}}(\bar{x})S({\bf g}_{2m+1}^{*})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}S({\bf g}_{2m+1})F_{0}^{(2m)}(y).$ (60)
Moreover,
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $-F_{1}^{(2m-1)^{*}}(\bar{x})S(\widetilde{\bf g}_{2m+1}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S(\widetilde{\bf g}_{2m+1})F_{1}^{(2m-1)}(y),$ (61)
where
$\widetilde{\bf g}_{2m+1}(z):=A_{0}z^{m-1}+A_{2}z^{m-2}+\ldots+A_{2m-2}.$ (62)
Proof.
We prove (58). From (41), (45) and (50), we have
${\mathscr{G}}_{2m}^{(1)}(x,y)=\widehat{\bf h}_{2m}^{(11)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(11)}(y).$ (63)
We examine $\widehat{\bf h}_{2m}^{(11)}(y)$. By using (54) and (55), we have
$\widehat{\bf h}_{2m}^{(11)}(y)=$ $\begin{pmatrix}A_{2m-2}+\ldots+A_{2}y^{2m-4}+I_{q}y^{2m-2}\\ A_{2m-4}+\ldots+A_{2}y^{2m-6}+I_{q}y^{2m-4}\\ \ldots\\ I_{q}\end{pmatrix}$
$=$ $\begin{pmatrix}A_{2m-2}&\ldots&A_{2}&I_{q}\\ A_{2m-4}&\ldots&I_{q}&\\ \ldots\\ I_{q}\end{pmatrix}\begin{pmatrix}I_{q}\\ y^{2}I_{q}\\ \ldots\\ y^{2m-2}I_{q}\end{pmatrix}$
$=$ $S({\bf h}_{2m})\left(I_{q},\bar{y}^{2}I_{q},\ldots,\bar{y}^{2m-2}I_{q}\right)^{*}.$ (64)
By substituting (64) for (63), we obtain (58). Similarly, by using
${\mathscr{G}}_{2m}^{(2)}(x,y)=$ $\widehat{\bf h}_{2m}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}\widehat{\bf h}_{2m}^{(21)}(y),$ (65)
${\mathscr{G}}_{2m+1}^{(1)}(x,y)=$ $\widehat{\bf g}_{2m+1}^{(11)^{*}}(\bar{x})\widetilde{J}_{m}H_{1,m-1}\widetilde{J}_{m}\,\widehat{\bf g}_{2m+1}^{(11)}(y),$ (66)
${\mathscr{G}}_{2m+1}^{(2)}(x,y)=$ $\widehat{\bf g}_{2m+1}^{(21)^{*}}(\bar{x})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}\widehat{\bf g}_{2m+1}^{(21)}(y),$ (67)
one can prove (59), (60) and (61). Equalities (65), (66) and (67) readily follow from (51), (52), (53) and (46), as well as from (47) and (48). Additionally, one also uses (42), (43), (44), (56) and (57). ∎
In the following corollary, we represent ${\mathscr{F}}_{n}$ as a matrix quadratic form regarding the block Hankel matrices $H_{1,j}$ and $H_{2,j}$, as well as the symmetrizer matrices $S({\bf h}_{2m})$, $S({\bf g}_{2m+1})$ and $S(\widetilde{\bf g}_{2m+1})$.
Corollary 3.5.
Assume the same conditions and notations as in Theorem 3.4. Let $\widetilde{J}_{m}$ be as in (49). Moreover, let
${\mathscr{J}}_{r}^{(\ell)}:=\begin{pmatrix}\widetilde{J}_{\ell}&0\\ 0&\widetilde{J}_{r}\end{pmatrix}$ (68)
for all $r,\,\ell>1$. Furthermore, for $k=0,1$ let
${\mathscr{W}}_{2m}^{(k)}:=$ $\begin{pmatrix}S({\bf h}_{2m}^{*})&0\\ 0&S({\bf h}_{2m}^{*})\end{pmatrix}{\mathscr{J}}_{m-1}^{(m-1)}\begin{pmatrix}H_{2,m-1}&0\\ 0&(-1)^{k}H_{1,m-1}\end{pmatrix}{\mathscr{J}}_{m-1}^{(m-1)}$
$\cdot\begin{pmatrix}S({\bf h}_{2m})&0\\ 0&S({\bf h}_{2m})\end{pmatrix}$, (69)
${\mathscr{W}}_{2m+1}^{(k)}:=$ $\begin{pmatrix}S({\bf g}_{2m+1}^{*})&0\\ 0&S(\widetilde{\bf g}_{2m+1}^{*})\end{pmatrix}{\mathscr{J}}_{m-1}^{(m)}\begin{pmatrix}H_{1,m}&0\\ 0&(-1)^{k}H_{2,m-1}\end{pmatrix}{\mathscr{J}}_{m-1}^{(m)}$
$\cdot\begin{pmatrix}S({\bf g}_{2m+1})&0\\ 0&S(\widetilde{\bf g}_{2m+1})\end{pmatrix}$. (70)
With ${\mathscr{F}}_{n}$ be as in (36), the following equalities hold:
${\mathscr{F}}_{2m}(x,y)=$ $2\begin{pmatrix}F_{0}^{(2m-2)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{W}}_{2m}^{(1)}\begin{pmatrix}F_{0}^{(2m-2)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}$, (71)
and
${\mathscr{F}}_{2m+1}(x,y)=$ $2\begin{pmatrix}F_{0}^{(2m)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{W}}_{2m+1}^{(1)}\begin{pmatrix}F_{0}^{(2m)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}$. (72)
Proof.
Equality (71) directly follows (39), (58), (59) and (69). In a similar manner, Equality (72) can readily be derived from (39), (60), (61) and (70). ∎
4 Hurwitness of Hurwitz-type matrix polynomials
In this section, using the explicit form of the Bezoutian matrix (71) and (72), we verify that an HTM polynomial ${\bf f}_{n}$ is indeed a Hurwitz matrix polynomial.
In [52], the Hurwitz property of HTM polynomials was studied. However, the proof of [52, Theorem 3.5] concerning the inertia of the matrix polynomial ${\bf f}_{n}$ for $n=2m$ is neither explicit nor complete, and a proof for the case $n=2m+1$ is not provided. See Remark 6.2.
Recall that the inertia of matrix polynomials, denoted by $\gamma({\bf f}_{n})$, provides the number of eigenvalues of ${\bf f}_{n}$ with positive, negative and real zero parts, respectively; see [43, page 410].
We rewrite Corollary 3.5 for the pair $(ix,iy)$ instead of $(x,y)$.
Corollary 4.1.
Assume the same conditions and notations as in Theorem 3.4. Moreover, let ${\mathscr{F}}_{n}$, ${\mathscr{J}}_{k}^{\ell}$, ${\mathscr{W}}_{2m}^{(0)}$ and ${\mathscr{W}}_{2m+1}^{(0)}$ be as in (36), (68), (69) and (70), respectively. Thus, the following equalities hold:
${\mathscr{F}}_{2m}(ix,iy)=$ $2\begin{pmatrix}F_{0}^{(2m-2)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}\begin{pmatrix}F_{0}^{(2m-2)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}$ (73)
and
${\mathscr{F}}_{2m+1}(ix,iy)=$ $2\begin{pmatrix}F_{0}^{(2m)}(\bar{x})\\ F_{1}^{(2m-1)}(\bar{x})\end{pmatrix}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}\begin{pmatrix}F_{0}^{(2m)}(y)\\ F_{1}^{(2m-1)}(y)\end{pmatrix}$. (74)
Proof.
Equalities (73) and (74) readily follow from (71) and (72), respectively. ∎
Let
$V_{[0,2m-1]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-2}I_{q},\bar{x}^{2m-1}I_{q}\right)^{\mathsf{*}},$
$V_{[0,2m]}(x):=$ $\left(I_{q},\bar{x}I_{q},\bar{x}^{2}I_{q},\ldots,\bar{x}^{2m-1}I_{q},\bar{x}^{2m}I_{q}\right)^{\mathsf{*}}.$
Furthermore, let ${\mathscr{T}}_{2m}$ (resp. ${\mathscr{T}}_{2m+1}$) be elementary block matrices such that
$\begin{pmatrix}F_{0}^{(2m-2)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}V_{[0,2m-1]}(x)$ (75)
and
$\begin{pmatrix}F_{0}^{(2m)}(x)\\ F_{1}^{(2m-1)}(x)\end{pmatrix}={\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(x).$ (76)
The following remark readily follows.
Remark 4.2.
Using (75) and (76), equalities (73) and (74) can be written as follows:
${\mathscr{F}}_{2m}(ix,iy)=$ $2V_{[0,2m-1]}^{*}(\bar{x}){\mathscr{T}}_{2m}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}V_{[0,2m-1]}(y),$ (77)
${\mathscr{F}}_{2m+1}(ix,iy)=$ $2V_{[0,2m]}^{*}(\bar{x}){\mathscr{T}}_{2m+1}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}V_{[0,2m]}(y).$ (78)
We now state the main result of the present work.
Theorem 4.3.
Each HTM polynomial is a Hurwitz matrix polynomial.
Proof.
We use Corollary 2.6 with
$L(\lambda)={\bf f}_{n}(i\lambda)\quad\mbox{and}\quad L_{1}(\lambda)={\bf f}_{n}(-i\lambda)$ (79)
and Equalities (29) , (30), (36), (37), (38) and (39). For $n=2m$, by using (77), we have
${\displaystyle\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})}=2{\mathscr{T}}_{2m}^{*}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{W}}_{2m}^{(0)}{\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}.$ (80)
By employing (78) for $n=2m+1$, we have
${\displaystyle\frac{1}{i}B_{L_{1}^{*},L^{*}}(L,L_{1})}=2{\mathscr{T}}_{2m+1}^{*}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{W}}_{2m+1}^{(0)}{\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}.$ (81)
The right-hand sides of (80) and (81) are positive definite matrices because
$\begin{pmatrix}H_{2,m-1}&0\\ 0&H_{1,m-1}\end{pmatrix}\quad\mbox{and}\quad\begin{pmatrix}H_{1,m}&0\\ 0&H_{2,m-1}\end{pmatrix}$
are positive definite, and both ${\mathscr{J}}_{m-1}^{(m-1)}{\mathscr{T}}_{2m}$ and ${\mathscr{J}}_{m-1}^{(m)}{\mathscr{T}}_{2m+1}$ are invertible matrices
From Corollary 2.6, the spectrum of HTM polynomials ${\bf f}_{2m}(i\lambda)$ and ${\bf f}_{2m+1}(i\lambda)$ belongs to the upper half-plane of $\mathbb{C}$}. Consequently, the spectrum of ${\bf f}_{2m}(\lambda)$ and ${\bf f}_{2m+1}(\lambda)$ lies in the left half-plane of $\mathbb{C}$}. Thus, polynomials ${\bf f}_{2m}(\lambda)$ and ${\bf f}_{2m+1}(\lambda)$ are Hurwitz matrix polynomials. ∎
Remark 4.4.
By employing (31), note that $\gamma_{-}({\bf f}_{n}(i\lambda))=\gamma_{0}({\bf f}_{n}(i\lambda))=0$ and $\gamma({\bf f}_{n}(i\lambda))=\gamma_{+}({\bf f}_{n}(i\lambda))$.
Remark 4.5.
A $q\times q$ HTM polynomial ${\bf f}_{2}$ of degree $2$ is a Hurwitz matrix polynomial; it admits the representation
${\bf f}_{2}(z)=I_{q}z^{2}+s_{0}z+s_{0}^{-1}s_{1}.$
Remark 4.6.
[…]
For $n=2m+1$ (resp. $n=2m$), the points of discontinuity of the distribution function $\widetilde{\sigma}_{2m,\max}$ (resp. $\widetilde{\sigma}_{2m,\min}$) that correspond to the measure $\sigma_{2m,\max}$ (resp. $\sigma_{2m,\min}$) are given by the roots of the polynomial $\det(P_{1,m}^{*}(-\bar{z}))$ (resp. $\det(zP_{2,m}^{*}(-\bar{z}))$); see (27) and (28).
On the other hand, the polynomials (25) and (26), which are constructed using the moments $s_{j}$, can be obtained from an absolutely continuous distribution on $0,\infty)$ that corresponds to a positive measure on $0,\infty)$, as shown in the next example.
Example 4.7.
Let $\widetilde{\sigma}(x)=\begin{pmatrix}4-2e^{-x/2}&ie^{-x}\\ -ie^{-x}&2-e^{-x}\\ \end{pmatrix}$ be defined on $0,\infty)$. We construct the moments $s_{0}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix}$, $s_{1}=\begin{pmatrix}4&-i\\ i&1\end{pmatrix}$, $s_{2}=\begin{pmatrix}16&-2i\\ 2i&2\end{pmatrix}$, $s_{3}=\begin{pmatrix}96&-6i\\ 6i&6\end{pmatrix}$, $s_{4}=\begin{pmatrix}768&-24i\\ 24i&24\end{pmatrix}$ of $\widetilde{\sigma}(x)$. The sequence $(s_{j})_{j=0}^{4}$ is a Stieltjes positive definite sequence. By using these moments and Definition 2.4, we derive $P_{2,2}$ and $Q_{2,2}$. From (24), we obtain the HMT polynomial
${\bf f}_{5}(z)=z^{5}I_{2}+A_{1}z^{4}+A_{2}z^{3}+A_{3}z^{2}+A_{4}z+A_{5},$ (82)
where $A_{1}=\begin{pmatrix}2&-i\\ i&1\end{pmatrix},A_{2}=\begin{pmatrix}\frac{486}{37}&0\\ -\frac{264}{37}i&6\end{pmatrix},A_{3}=\begin{pmatrix}\frac{560}{37}&-5i\\ 5i&5\end{pmatrix}$,
$A_{4}=\begin{pmatrix}\frac{1158}{37}&0\\ -\frac{936}{37}i&6\\ \end{pmatrix}$ and $A_{5}=\begin{pmatrix}\frac{292}{37}&-2i\\ 2i&2\\ \end{pmatrix}$. Therefore, by Theorem 4.3, the polynomial ${\bf f}_{5}$ is a Hurwitz polynomial.
Next, we paraphrase [5, Theorem 7.10].
Theorem 4.8.
Let $n$ be a natural number with $n\geq 2$, and let ${\bf f}_{n}$ be a matrix-valued polynomial. Let ${\bf h}_{n}$ (resp. ${\bf g}_{n}$) denote the ${\bf h}_{n}$-part (resp. ${\bf g}_{n}$-part) of ${\bf f}_{n}$. Furthermore, let $n=2m$ (resp. $n=2m+1$) with $m\geq 1$, and let $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) be constructed from the Hermitian Markov parameters of the polynomial ${\bf f}_{n}$. Then the matrix-valued polynomial ${\bf f}_{n}$ is an HTM polynomial if and only if the block Hankel matrices $H_{1,m-1}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m-1}$) are positive definite.
Proposition 4.9.
Let ${\bf f}_{n}$ be an HTM polynomial. Moreover, let ${\bf h}_{n}$ and ${\bf g}_{n}$ be as in (2), (5) and (8), respectively. Furthermore, let $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ and $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ be Bezoutians associated with the quadruple $(x{\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},x{\bf g}_{n})$ and $({\bf g}_{n}^{*},{\bf h}_{n},{\bf h}_{n}^{*},{\bf g}_{n})$, respectively. Thus,
a) the matrix $B_{x{\bf g}_{n}^{*},{\bf h}_{n}^{*}}(x{\bf g}_{n},{\bf h}_{n})$ is positive definite, and
b) the matrix $B_{{\bf g}_{n}^{*},{\bf h}_{n}^{*}}({\bf g}_{n},{\bf h}_{n})$ is negative definite.
Proof.
Let $\widetilde{J}_{m}$, $\widetilde{\bf g}_{2m+1}$, $S({\bf h}_{2m})$, $S({\bf g}_{2m+1})$ and $S(\widetilde{\bf g}_{2m+1})$ be as in (49), (62) and (55). For $n=2m$, by using (37), (38), (58) and (59), we have
$B_{x{\bf g}_{2m}^{*},{\bf h}_{2m}^{*}}(x{\bf g}_{2m},{\bf h}_{2m})=$ $S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m}),$
$B_{{\bf g}_{2m}^{*},{\bf h}_{2m}^{*}}({\bf g}_{2m},{\bf h}_{2m})=$ $-S({\bf h}_{2m}^{*})\widetilde{J}_{m-1}H_{1,m-1}\widetilde{J}_{m-1}S({\bf h}_{2m}).$
By (37), (38), (60) and (61), for $n=2m+1$, we obtain
$B_{x{\bf g}_{2m+1}^{*},{\bf h}_{2m+1}^{*}}(x{\bf g}_{2m+1},{\bf h}_{2m+1})=$ $S({\bf g}_{2m+1}^{*})\widetilde{J}_{m}H_{1,m}\widetilde{J}_{m}S({\bf g}_{2m+1})$
$B_{{\bf g}_{2m+1}^{*},{\bf h}_{2m+1}^{*}}({\bf g}_{2m+1},{\bf h}_{2m+1})=$ $-S(\widetilde{\bf g}_{2m+1}^{*})\widetilde{J}_{m-1}H_{2,m-1}\widetilde{J}_{m-1}S(\widetilde{\bf g}_{2m+1}).$
From Theorem 4.8, see that assertion a) and b) are valid. ∎
Remark 4.10.
In [42, Theorem 2], it is proven that a scalar polynomial $f_{n}$ is Hurwitz if and only if $B_{xg_{n}^{*},h_{n}^{*}}(xg_{n},h_{n})$ is negative definite and $B_{g_{n}^{*},h_{n}^{*}}(g_{n},h_{n})$ is positive definite. Note that the Markov parameters related to [42, Theorem 2] have the opposite sign compared with those associated with parts (a) and (b) of Proposition 4.9.
5 Completing a matrix polynomial to an HTM polynomial
As indicated in [9], not every Hurwitz matrix polynomial is an HTM polynomial. In this section, we propose a procedure to complete a given matrix polynomial into an HTM polynomial.
In [9], the polynomial
$P_{3}(z)=I_{2}z^{3}+A_{1}z^{2}+A_{2}z+A_{3}$ (83)
with
$A_{1}=\begin{pmatrix}\frac{2187}{109}&\frac{1206}{109}\\ 0&9\end{pmatrix},\quad A_{2}=\begin{pmatrix}\frac{10662}{109}&\frac{8700}{109}\\ 0&18\end{pmatrix},\quad A_{3}=\begin{pmatrix}\frac{11178}{109}&\frac{10524}{109}\\ 0&6\end{pmatrix}$ (84)
was suggested as a polynomial that is Hurwitz but not an HTM polynomial. The rational matrix function $\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}$ does not satisfy property (35). Moreover, as indicated in [9], the $2\times 2$ matrix coefficients $c_{j}$ of the expansion
$\frac{1}{z}\left(A_{1}z+A_{3}\right)\left(I_{2}z+A_{2}\right)^{-1}=-\frac{c_{0}}{z}-\frac{c_{1}}{z^{2}}-\frac{c_{2}}{z^{3}}$-$\cdots$
are not Hermitian matrices. Consequently, the polynomial $P_{3}$ cannot be investigated in the framework of HTM polynomials.
The following result allows one to determine whether a matrix polynomial $P_{n}$ that is not an HTM polynomial is in fact a Hurwitz polynomial by “completing,” mainly by perturbing $P_{n}$ with another polynomial $Q_{n-1}$. This is done by constructing a new HTM polynomial of degree $2n$.
Theorem 5.1.
Let $P_{n}(z)$ be $q\times q$ a monic polynomial of degree $n$. If there exists a $q\times q$ polynomial $Q_{n-1}(z)$ of degree n1n-1 such that
${\bf f}_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ (85)
is an HTM polynomial, then the polynomial $P_{n}(z)$ is a Hurwitz polynomial.
Proof.
From (24) and (25), there are polynomials $P_{1,n}$ y $Q_{1,n}$ such that $P_{n}(z)=(-1)^{n}P_{1,n}^{*}(-\bar{z})$ $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$ and $P_{1,n}(z)$ is an orthogonal polynomial on $0,+\infty)$ with respect to a certain positive measure $\sigma$ corresponding to the Stieltjes transform $\frac{Q_{1,n}^{*}(\bar{z})}{P_{1,n}^{*}(\bar{z})}$. By [22, Theorem 1], the roots of the polynomial $\det P_{1,n}(z)$ lie in $0,+\infty)$. Consequently, the roots of polynomial $\det P_{n}(z)=\det((-1)^{n}P_{1,n}^{*}(-\bar{z}))$ lie in $(-\infty,0]$. ∎
Remark 5.2.
Let ${\bf f}_{n}$ be a monic matrix polynomial and ${\bf f}_{n}^{\intercal}$ be its transpose. Since $\det{\bf f}_{n}=\det{\bf f}_{n}^{\intercal}$, Theorem 5.1 is valid for $P_{n}^{\intercal}$ and $Q_{n-1}^{\intercal}$.
For finding polynomial $Q_{n-1}$ as in Theorem 5.1, one can use the following algorithm.
Algorithm for finding $Q_{n-1}$ from $P_{n}$
Step 1. Let $P_{1,n}(z)=(-1)^{n}P_{n}^{*}(-\bar{z})$ (resp. $P_{1,n}(z)=(-1)^{n}P_{n}^{\intercal^{*}}(-\bar{z})$). By using (22), compute $Q_{1,n}(z)$ in terms of the coefficients of $P_{n}(z)$ (resp. $P_{n}^{\intercal}(z)$) and the moments or Markov parameters $s_{j}$, associated with some positive measure $\sigma$ on $0,+\infty)$.
Step 2. Verify that $\frac{P_{1,n}^{*}(\bar{z})}{Q_{1,n}^{*}(\bar{z})}$ satisfies Equality (35). If Equality (35) is not satisfied, we cannot apply this algorithm.
Step 3. By using the rational matrix function $\frac{P_{1,n}^{*}(\bar{z})}{Q_{1,n}^{*}(\bar{z})}$, calculate the remaining Markov parameters from the set $(s_{j})_{j=0}^{2n-1}$.
Step 4. If sequence $(s_{j})_{j=0}^{2n-1}$ is a positive definite Stieltjes sequence, meaning that the corresponding Hankel matrices $H_{1,n-1}$ and $H_{2,n-1}$ are positive definite matrices, then $Q_{n-1}(z)=(-1)^{n+1}Q_{1,n}^{*}(-\bar{z})$.
Example 5.3.
We apply the completion procedure to the matrix polynomial (83). Let the superscript $\mathsf{T}$ denote the transpose of a matrix. For $P_{3}^{\intercal}(z)$, we use the above algorithm together with Remark 5.2. With moments $s_{0}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix}$, $s_{1}=\begin{pmatrix}4b&-b\\ -b&b\end{pmatrix}$ and $s_{2}=\begin{pmatrix}16b&-2b\\ -2b&2b\end{pmatrix}$, we calculate $Q_{1,3}(z)$; therefore, we have
$Q_{1,3}(z)=s_{0}z^{2}+z(s_{1}-A_{1}^{\intercal^{*}}s_{0})+s_{2}-A_{1}^{\intercal^{*}}s_{1}+A_{2}^{\intercal^{*}}s_{0}.$
From the Laurent expansion of $\frac{P_{1,3}^{*}(\bar{z})}{Q_{1,3}^{*}(\bar{z})}$, we calculate the moments
$s_{3}=\begin{pmatrix}96b&-6b\\ -6b&6b\end{pmatrix},s_{4}=\begin{pmatrix}768b&-24b\\ -24b&24b\end{pmatrix},s_{5}=\begin{pmatrix}7680b&-120b\\ -120b&120b\end{pmatrix}$.
The block Hankel matrices $H_{1,2}$ and $H_{2,2}$ are positive definite matrices for $b>0$. From (25), polynomial $Q_{2}(z)$ constructed from $P_{3}$ has the form $Q_{2}(z)=Q_{1,3}^{*}(-\bar{z})$. Thus, the resulting HTM polynomial ${\bf f}_{6}$ can be represented as in (85):
${\bf f}_{6}(z)=z^{6}I_{2}+B_{1}z^{5}+A_{1}z^{4}+B_{2}z^{4}+A_{2}z^{2}+B_{3}z+A_{3}.$ (86)
The coefficients $A_{1}$, $A_{2}$ and $A_{3}$ are given by (84), and the coefficients $B_{1}$, $B_{2}$ and $B_{3}$ are
$B_{1}=\begin{pmatrix}2b&-b\\ -b&b\end{pmatrix},B_{2}=\begin{pmatrix}\frac{2732b}{109}&-8b\\ -8b&8b\\ \end{pmatrix},B_{3}=\begin{pmatrix}\frac{6826b}{109}&-11b\\ -11b&11b\end{pmatrix}$.
By Theorem 4.3, matrix polynomial (86) is a Hurwitz matrix polynomial for $b>0$. Finally, by Theorem 5.1, polynomial $P_{3}$ as in (83) is a Hurwitz matrix polynomial.
In the scalar version, one can construct a sequence of scalar Hurwitz polynomials $(f_{j})_{j=1}^{\infty}$ using the sequence of scalar orthogonal polynomials on $0,+\infty)$ $(p_{1,j})_{j=1}^{\infty}$ and $(p_{2,j})_{j=1}^{\infty}$ as in Definition 2.4; see [11].
In the following lemma, we propose constructing a family of scalar Hurwitz polynomials involving the polynomials $p_{1,j}$, and $p_{2,j}$ and their derivatives.
Lemma 5.4.
For $q=1$, let $(p_{1,m}(z))$ (resp. $(p_{2,m}(z))$) be as in Definition 2.4, and let $\widetilde{p}_{1,m-1}(z):=\frac{d}{dz}p_{1,m}(z)$ (resp. $\widetilde{p}_{2,m-1}(z):=\frac{d}{dz}p_{2,m}(z)$). Thus, for $m\geq 2$, the polynomials
$f_{2m}(z)=(-1)^{m}(p_{1,m}(-z^{2})-z\widetilde{p}_{1,m-1}(-z^{2}))$ (87)
and
$f_{2m+1}(z)=(-1)^{m}(\widetilde{p}_{2,m}(-z^{2})+zp_{2,m}(-z^{2})).$ (88)
are scalar Hurwitz polynomials.
Proof.
The proof of (87) and (88) readily follows from the interlacing properties of the roots of the orthogonal polynomials $p_{k,m}(z)$ and $\widetilde{p}_{k,m-1}(z)$ for $k=1,2$, and from the fact that $a_{0}a_{1}>0$, where $a_{0}$ and $a_{1}$ are the leading coefficients of the polynomial $f_{n}(z)=a_{0}z^{n}+a_{1}z^{n-1}+\cdots+a_{n}$; see the Hermite–Biehler theorem [3], [33], [48, page 46]. ∎
Remark 5.5.
From polynomial (82), we observe that the coefficients of an HTM polynomial may be complex matrices, whereas in the scalar case the necessary condition for Hurwitzness requires all coefficients of a Hurwitz polynomial to be positive. On the other hand, from (86) and (82), we see that the determinants of the coefficients of these polynomials are positive. Based on this observation, we propose the conjecture below.
Conjecture. The determinants of the coefficients $A_{j}$ of an HTM polynomial ${\bf f}_{n}$ are positive.
6 Additional remarks
In this section, to clarify the importance of the present work, we recall several relevant results from [52], paying particular attention to certain steps in the proofs that were left implicit. We emphasize that the authors of [52] (2021) were the first to propose that the HTM polynomials possess the Hurwitz property, while the concept of HTM polynomials itself was introduced earlier by the author of the present work in [5] (2015).
Lemma 6.1.
[52, Lemma 3.4] Let $N_{R}(z)$, $D_{R}(z)$, $N_{L}(z)$ and $D_{L}(z)\in\mathbb{C}[z]^{p\times p}$ be regular such that $\deg D_{R}>\deg N_{R}$, $\deg D_{L}>\deg N_{L}$ and, for all large enough $z\in\mathbb{C}$,
$(D_{L}(z))^{-1}N_{L}(z)=N_{R}(z)(D_{R}(z))^{-1}=\sum_{k=0}^{\infty}z^{-(k+1)}s_{k}.$
If $D_{L}(z)$ and $D_{R}(z)$ are written in the form
$D_{L}(z)=\sum_{k=0}^{m_{L}}D_{L,m_{L}-k}z^{k}\qquad\text{and}\qquad D_{R}(z)=\sum_{k=0}^{m_{R}}D_{R,m_{R}-k}z^{k},$
where $D_{L,k}\in\mathbb{C}^{p\times p}$ for $k=0,\ldots,m_{L}$ and $D_{R,k}\in\mathbb{C}^{p\times p}$ for $k=0,\ldots,m_{R}$, then
$\mathbf{B}_{D_{L},N_{L}}(D_{R},N_{R})=$ $\begin{pmatrix}D_{L,m_{L}-1}&\cdots&D_{L,0}\\ \vdots&\ddots&\vdots\\ D_{L,0}&\cdots&\end{pmatrix}\begin{pmatrix}s_{0}&\cdots&s_{m_{R}-1}\\ \vdots&\ddots&\vdots\\ s_{m_{L}-1}&\cdots&s_{m_{R}+m_{L}-2}\end{pmatrix}$
$\cdot\begin{pmatrix}D_{R,m_{R}-1}&\cdots&D_{R,0}\\ \vdots&\ddots&\vdots\\ D_{R,0}&\cdots&\end{pmatrix}$. (89)
Let $L_{1}$ and $L$ be as in (79). In [52], the Theorem 3.5 is regarding the inertia representation of matrix polynomials ${\bf f}_{n}$. For $n=2m$, in the proof of [52, Theorem 3.5] the [52, Equality (3.6)] in terms of our notations is the following
$B_{L_{1}^{*},L^{*}}(L,L_{1})=2iB_{{\bf h}_{2m}^{*}(-\bar{x}^{2}),x{\bf g}_{2m}^{*}(-\bar{x}^{2})}({\bf h}_{2m}(-x^{2}),-x{\bf g}_{2m}(-x^{2})).$ (90)
The authors of [52] use $2i\,B_{F_{e}^{\vee},\,-F_{0}^{\vee}}\bigl(\widehat{F}_{e},\,-\widehat{F}_{0}\bigr)$ in place of the right-hand side of (90). They claim that Lemma 6.1 implies that the matrix $2i\,B_{F_{e}^{\vee},\,-F_{0}^{\vee}}\bigl(\widehat{F}_{e},\,-\widehat{F}_{0}\bigr)$ is congruent to the block Hankel matrix
$\begin{pmatrix}s_{0}&0_{p}&s_{1}&\cdots&s_{m-1}&0_{p}\\ 0_{p}&s_{1}&0_{p}&\cdots&0_{p}&s_{m}\\ s_{1}&0_{p}&s_{2}&\cdots&s_{m}&0_{p}\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ s_{m-1}&0_{p}&s_{m}&\cdots&s_{2m-2}&0_{p}\\ 0_{p}&s_{m}&0_{p}&\cdots&0_{p}&s_{2m-1}\end{pmatrix}$, (91)
and that, after a suitable reordering of rows and columns in (91), it is further congruent to
$\begin{pmatrix}H_{1,m-1}&0\\ 0&H_{2,m-1}\end{pmatrix}$. (92)
Remark 6.2.
By comparing the Hankel matrix of the right-hand side of (89) and the matrix (91) It is not clear how from $2i\,B_{F_{e}^{\vee},\,-F_{0}^{\vee}}\bigl(\widehat{F}_{e},\,-\widehat{F}_{0}\bigr)$ one gets (91). Furthermore, it is not clear how from (91) one obtains (92).
In contrast, for $n=2m$, in the present work we derive the matrix $\left(\begin{smallmatrix}H_{2,m-1}&0\\ 0&H_{1,m-1}\end{smallmatrix}\right)$ associated with the quadratic form in (92) by employing the decomposition (39) and by computing explicitly the quadratic forms corresponding to (37) and (38), as shown in (71). Additionally, for $n=2m+1$, the proof of [52, Theorem 3.5] is not given.
In Theorem 4.10 of [52], a “stability criterion via Markov parameters” for matrix polynomials ${\bf f}_{n}$ in even and odd cases is suggested. We have the following remark regarding this.
Remark 6.3.
For $n=2m$, Theorem 4.10 of [52], is related to the HTM polynomials. Other matrix polynomials, for example, the Hurwitz polynomial (83), are not covered by [52, Theorem 4.10].
In this work, we do not discuss the case $n=2m+1$ of Theorem 4.10 of [52].
7 Conclusion
In this work we obtained an explicit representation of the Bezoutian associated with Hurwitz-type matrix polynomials. This representation clarifies the relation between Hurwitz-type matrix polynomials and classical Hurwitz matrix polynomials and provides a constructive tool for verifying Hurwitzness in the matricial setting. Our approach, based on the decomposition of Bezoutians and on the structure of truncated matricial Stieltjes moment problems, offers a natural framework for extending classical stability criteria to matrix polynomials. In addition, we proposed a method to enlarge the class of Hurwitz-type matrix polynomials by perturbing a polynomial that is not of Hurwitz type so that the new polynomial becomes Hurwitz type. These results contribute to a better understanding of the structure of Hurwitz-type matrix polynomials and their relation to Bezoutians, and they suggest further developments in matrix moment problems, continued fractions, and stability theory.